### v1

# Ruishu Liu, Yulin Cao,  Xiaojie Wang

Abstract
The present work is devoted to strong approximations of a generalized Aït-Sahalia model arising from mathematical finance. The numerical study of the considered model faces essential difficulties caused by a drift that blows up at the origin, highly nonlinear drift and diffusion coefficients and positivity-preserving requirement. In this paper, a novel explicit Euler-type scheme is proposed, which is easily implementable and able to preserve positivity of the original model unconditionally, i.e., for any time step-size $h>0$. A mean-square convergence rate of order $0.5$ is also obtained for the proposed scheme in both non-critical and general critical cases. Our work is motivated by the need to justify the multi-level Monte Carlo (MLMC) simulations for the underlying model, where the rate of mean-square convergence is required and the preservation of positivity is desirable particularly for large discretization time steps. To the best of our knowledge, this is the first paper to propose an unconditionally positivity preserving explicit scheme with order $1/2$ of mean-square convergence for the model. Numerical experiments are finally provided to confirm the theoretical findings.
Keywords: Aït-Sahalia model; Explicit scheme; Unconditionally positivity preserving; Strong convergence rates.
1 Introduction
The pricing of financial assets has been an important issue in mathematical finance over the past few decades, where stochastic differential equations (SDEs) are prominent models frequently used to capture the behaviors of financial systems. Notable examples include the Black-Scholes model [2], the Merton model [14] and the Cox-Ingersoll-Ross (CIR) model [5] in the modern financial theory. The economist and mathematician Aït-Sahalia conducted an empirical study in [1] and investigated several continuous-time models for interest rates. He tested the models by comparing the parameter density implied by the parameter models, and found that all the existing univariate linear drift models fail to explain the euro dollar well. As a consequence of this work, a nonlinear stochastic interest model which captures well of the dynamics of the spot rate was proposed and named as the Aït-Sahalia model, which is now widely used in finance and economics, and described by the following Itô-type SDE:
$\mathrm{d}X_{t}=\big{(}c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{2}\big{)}\,\text{d}t+c_{3}X_{t}^{\rho}\,\text{d}W_{t},\quad X_{0}=x_{0}>0.$ (1.1)
Here $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\rho>1$ and $\{W_{t}\}_{t\in[0,T]}$ stands for a standard Brownian motion defined on a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Aït-Sahalia [1] proved that the model (1.1) admits a unique global positive solution using the Feller explosion test. In 2009, Cheng [4] derived the finite moment bounds for the analytic solutions and proved that the Euler–Maruyama approximate solutions are convergent in probability. Later in 2011, Szpruch et al. [18] extended the model (1.1) to a more general version given by
$\mathrm{d}X_{t}=$ $(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\text{d}t+c_{3}X_{t}^{\rho}\text{d}W_{t},\quad X_{0}=x_{0}>0,\>\kappa,\rho>1,$ (1.2)
and showed that the generalized model admits a unique global positive solution with uniformly in time bounded $p$-th moments ($p\in\mathbb{R}$) in the non-critical case $\kappa+1>2\rho$. Moreover, they applied the classical backward Euler-Maruyama (BEM) scheme to approximate the model and proved the positivity-preserving approximations to be strongly convergent under the condition that $\kappa+1>2\rho$, but with no convergence rates revealed. Recently, the authors of [20] successfully recovered the mean-square convergence rate of order $1/2$ for the stochastic theta methods (STMs) with $\theta\in[\tfrac{1}{2},1]$ applied to the model (1.2) under the condition that $\kappa+1\geq 2\rho$, covering the critical case $\kappa+1=2\rho$. Also, we mention that Neuenkirch and Szpruch [15] proposed a kind of Lamperti-backward Euler method for scalar SDEs defined in a domain, including the Ait-Sahalia-type model (1.2), with a strong convergence rate of order one obtained. However, the computational costs of solving implicit algebraic equations rise as the parameter $\kappa$ increases. More recently, a truncated explicit Euler method was proposed in [7] to approximate (1.2), but with no convergence rate revealed. When we finish the draft of the present work, we are also aware of two closely related publications [6, 13]. The authors of [6] recovered a mean-square convergence rate of order nearly $\tfrac{1}{4}$ for the truncated Euler method only in the non-critical case $\kappa+1>2\rho$. The work [13] proposed an exponential tamed method based on the Lamperti transform and splitting, which is strongly convergent with order one, but not strictly positivity preserving.
In the present article we aim to design novel explicit positivity preserving schemes for approximations of the Aït-Sahalia model (1.2) with mean-square convergence rate of order $\tfrac{1}{2}$. More precisely, on a uniform mesh with a uniform step-size $h=\tfrac{T}{N},\,N\in\mathbb{N}$ we propose the following time-stepping scheme:
$Y_{n+1}=Y_{n}+c_{-1}Y_{n+1}^{-1}h+\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}h+g_{h}(Y_{n})\Delta W_{n},\quad Y_{0}=X_{0},$ (1.3)
where $f_{h}$ and $g_{h}$ are some explicit modifications of $f(x):=-c_{2}x^{\kappa}$ and $g(x):=c_{3}x^{\rho}$ in the SDE model, respectively (see Section 3 for details). Clearly, the implicitness only appears in the term $c_{-1}Y_{n+1}^{-1}h$, which is the key element to preserve the positivity of the original model (see Lemma 3.1). On the contrary, the remaining terms are explicitly handled. It is not difficult to see that, given the former step $Y_{n}$, the current step $Y_{n+1}$ can be explicitly known, as the unique positive root of a quadratic equation (see (3.5) below for the explicit solution). In this way, the computational costs are significantly reduced compared with the implicit schemes (BEM or STMs) in existing publications [18, 20], where implicit algebraic equations are required to be numerically solved for every step. Furthermore, the proposed explicit Euler-type scheme is positivity-preserving for any time step-size $h>0$, which is particularly desirable for the multi-level Monte Carlo (MLMC) simulations [9] with large discretization time steps.
To justify the MLMC simulations, the mean-square convergence rate of the approximations is also necessary [9, Theorem 3.1]. Recall that the analysis of convergence rates in [20] for the implicit schemes does not rely on a priori moment estimates of numerical approximations. Nevertheless, due to the presence of the explicit treatment in the drift term, the analysis of convergence rates for the proposed semi-implicit schemes relies on finite moment bounds of numerical approximations:
$\sup_{N\in\mathbb{N}}\sup_{0\leq n\leq N}\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}<\infty,\quad p\geq 2.$ (1.4)
Deriving such $p$-th moment integrability of the numerical approximations is, however, not an easy task and some non-standard arguments are exploited (see the proof of Lemma 3.6), to overcome essential difficulties caused by a drift that blows up at the origin, highly nonlinear diffusion coefficients and a mixture of implicitness and explicitness in the drift part of the scheme. Equipped with the moment bound (1.4), one is able to identify the mean-square convergence rate of order $0.5$ for the proposed scheme (3.1) applied to the model (1.2) (see Theorem 4.1):
$\sup_{0\leq n\leq N}\mathbb{E}\big{[}\big{|}X_{t_{n}}-Y_{n}\big{|}^{2}\big{]}\leq Ch$ (1.5)
in both non-critical ($\kappa+1>2\rho$) and general critical ($\kappa+1=2\rho$) cases, where the constant $C$ is independent of the time step-size $h$. To the best of our knowledge, this is the first paper to devise an unconditionally positivity preserving explicit scheme with order $\tfrac{1}{2}$ of mean-square convergence for the general Aït-Sahalia model.
The rest of this paper is organized as follows. In the next section, basic notations and properties of the original model are presented. In Section 3 the newly proposed scheme is introduced and proved to be positivity preserving and moment bounded. The mean-square error bounds for the proposed scheme is carefully analyzed in Section 4, with order one-half of convergence revealed. Numerical experiments are shown in Section 5 to verify our theoretical results and a short conclusion is made in Section 6.
2 The generalized Aït-Sahalia model
[…]
$\|\xi\|_{\mathcal{L}^{p}(\Omega;\mathbb{R})}:=\big{(}\mathbb{E}\big{[}|\xi|^{p}\big{]}\big{)}^{1/p}<\infty.$ (2.1)
In this work we focus on the generalized Aït-Sahalia model taking the form as follows:
$\text{d}X_{t}=(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\text{d}t+c_{3}X_{t}^{\rho}\text{d}W_{t},\quad X_{0}=x_{0},\quad t>0,$ (2.2)
where $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$ and $\kappa+1\geq 2\rho$. For the particular case $\kappa=2$, the underlying model (2.2) reduces into (1.1), originally proposed by Aït-Sahalia [1]. The well-posedness of the generalized model (2.2) has been established in the following theorem quoted from [18, Theorem 2.1].
Theorem 2.1.
Let $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$. Given any initial value $X_{0}>0$, there exists a unique, positive global solution $X_{t}$ to the SDE (2.2) on $t\geq 0$.
Furthermore, we recall the moment integrability properties of the solution to (2.2) as follows (see [18, Lemma 2.1] and [20, Lemma 4.6] for details).
Lemma 2.2.
Let conditions in Theorem 2.1 be satisfied with $\kappa+1>2\rho$. For any $p\geq 2$, it holds
[…]
For the general critical case $\kappa+1=2\rho$, we have slightly different assertions, as stated in the following lemma (cf. [20, Lemma 4.6]).
Lemma 2.3.
Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $12\leq p_{1}\leq\tfrac{2c_{2}}{c_{3}^{2}}+$ 1 and for any $p_{2}\geq 2$ we have
[…]
Based on the above moment integrability properties, one can also obtain the Hölder continuity of the solutions, as stated in the following lemmas quoted directly from [20].
Lemma 2.4.
[20, Lemma 4.4] Let conditions in Theorem 2.1 be satisfied with $\kappa+1>2\rho$. Then it holds that, for any $p\geq 1$ and $t,s\in[0,T]$,
$\|X_{t}-X_{s}\|_{L^{p}(\Omega;\mathbb{R})}$ $\leq C|t-s|^{\tfrac{1}{2}},$ (2.5)
$\|X_{t}^{-1}-X_{s}^{-1}\|_{L^{p}(\Omega;\mathbb{R})}$ $\leq C|t-s|^{\tfrac{1}{2}}.$ (2.6)
Lemma 2.5.
[20, Lemma 4.7] Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $t,s\in[0,T]$ it holds that
$\|X_{t}-X_{s}\|_{L^{q_{1}}(\Omega;\mathbb{R})}$ $\leq C|t-s|^{\tfrac{1}{2}},\quad 2\leq q_{1}\leq\tfrac{1}{\kappa}\big{(}\tfrac{2c_{2}}{c_{3}^{2}}+1\big{)},$ (2.7)
$\|X_{t}^{-1}-X_{s}^{-1}\|_{L^{q_{2}}(\Omega;\mathbb{R})}$ $\leq C|t-s|^{\tfrac{1}{2}},\quad 2\leq q_{2}<\tfrac{1}{\kappa}\big{(}\tfrac{2c_{2}}{c_{3}^{2}}+1\big{)}.$ (2.8)
Here and below, for notation simplicity the letter $C$ is slightly abused to denote a generic positive constant, which is independent of the time step-size and may vary for each appearance.
As a straightforward consequence of the above lemmas, one can use the Hölder inequality to show the following estimates (consult [20, (4.21), (4.24)]).
Lemma 2.6.
Let conditions in Theorem 2.1 be satisfied with $\kappa+1\geq 2\rho$. If one of the following conditions stands:
  1. (1)
    $\kappa+1>2\rho$,
  2. (2)
    $\kappa+1=2\rho$, $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$,
then for any $t,s\in[0,T]$ it holds that
$\|f(X_{t})-f(X_{s})\|_{L^{2}(\Omega;\mathbb{R})}\leq C|t-s|^{\frac{1}{2}},\quad\|g(X_{t})-g(X_{s})\|_{L^{2}(\Omega;\mathbb{R})}\leq C|t-s|^{\frac{1}{2}},$ (2.9)
where we denote $f(x):=-c_{2}x^{\kappa}$ and $g(x):=c_{3}x^{\rho}$ for short.
Armed with the above properties of the underlying model, one can turn to the numerical approximations and analyze the resulting mean-square errors in the forthcoming sections.
3 The explicit positivity-preserving Euler-type scheme
In this section we aim to propose and analyze new strong approximation schemes for the generalized Aït-Sahalia model (2.2). For $T\in(0,+\infty)$ and $N\in\mathbb{N}$, we construct a uniform mesh $\{t_{n}=nh\}_{n=0}^{N}$ on the time interval $[0,T]$ with a uniform step-size $h=\tfrac{T}{N}$. Based on the uniform mesh, we propose an explicit Euler-type method for (2.2) as follows:
$Y_{n+1}=Y_{n}+c_{-1}Y_{n+1}^{-1}h+\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}h+g_{h}(Y_{n})\Delta W_{n},\quad Y_{0}=X_{0},$ (3.1)
where $\Delta W_{n}:=W_{t_{n+1}}-W_{t_{n}}$ for $n=0,1,...,N-1$ and $f_{h},g_{h}:(0,+\infty)\rightarrow\mathbb{R}$ are some modifications of
$f(x):=-c_{2}x^{\kappa}\quad\text{and}\quad g(x):=c_{3}x^{\rho},$ (3.2)
satisfying certain assumptions specified subsequently (cf. Assumptions 3.2-3.4). At first, we show that the proposed scheme preserves the positivity of the original model unconditionally.
Lemma 3.1 (Unconditional positivity-preserving).
Let $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$. Given any initial value $Y_{0}=X_{0}>0$, the scheme (3.1) admits unique, positive numerical solutions $\{Y_{n}\}_{n\in\mathbb{N}}$ for any stepsize $h=\tfrac{T}{N}>0$. In other words, the scheme (3.1) is positivity preserving unconditionally.
Proof: For $x>0$ we introduce the function
$G(x):=x-c_{-1}hx^{-1}.$ (3.3)
Evidently, the well-posedness of the scheme (3.1) in the positive domain $(0,\infty)$ is to find a unique positive solution of the implicit equation $G(x)=c$ for any $c\in\mathbb{R}$. On the one hand, we observe that $c_{-1}h>0$ and thus
$\lim_{x\rightarrow 0+}G(x)=-\infty,\quad\lim_{x\rightarrow\infty}G(x)=+\infty.$ (3.4)
On the other hand, note that the function $G$ is monotonically increasing in the domain $(0,\infty)$ since $G^{\prime}(x)=1+c_{-1}hx^{-2}>0$. As a direct consequence, for any $c\in\mathbb{R}$, $G(x)=c$ admits a unique root in $(0,+\infty)$ and this validates the desired assertion. ∎
Indeed, the unique positive solution determined by the proposed scheme (3.1) is given by
$Y_{n+1}=\frac{Y_{n}+\vartheta_{n}h+g_{h}(Y_{n})\Delta W_{n}+\sqrt{(Y_{n}+\vartheta_{n}h+g_{h}(Y_{n})\Delta W_{n})^{2}+4c_{-1}h}}{2},$ (3.5)
where for short we denote
$\vartheta_{n}:=-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n}).$
To prove the bounded moments of the numerical approximations, we formulate the following assumptions.
Assumption 3.2.
For any $x\in(0,+\infty)$, it holds that
$\big{|}f_{h}(x)\big{|}\leq Ch^{-\frac{1}{2}},\quad\big{|}g_{h}(x)\big{|}\leq Ch^{-\frac{1}{2}}.$ (3.6)
Assumption 3.3.
Let $\kappa+1\geq 2\rho$. There exist a positive constant $L(\gamma)$, independent of the stepsize $h>0$, such that
$\langle x,f_{h}(x)\rangle+\gamma|g_{h}(x)|^{2}\leq L<\infty,\quad\forall x\in(0,+\infty),$ (3.7)
where $\gamma>0$ is assumed to be arbitrarily large for the case $\kappa+1>2\rho$ and $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ for the critical case $\kappa+1=2\rho$.
Assumption 3.4.
There exists some positive constant $m_{1},m_{2}\leq 2\gamma+1$ such that
$\big{|}f(x)-f_{h}(x)\big{|}^{2}\leq Ch|x|^{m_{1}},\quad\big{|}g(x)-g_{h}(x)\big{|}^{2}\leq Ch|x|^{m_{2}},\quad\forall x\in(0,+\infty)$ (3.8)
where $f,g$ are defined by (3.2) and $\gamma$ comes from Assumption 3.3.
Next we show an example of modifications which meets all the above assumptions.
Example 3.5.
For $\kappa+1\geq 2\rho$, let us consider an example of a modification of $f,g$ as follows:
$f_{h}(x):=-\tfrac{c_{2}x^{\kappa}}{1+h^{\alpha}x^{2\kappa\alpha}},\quad g_{h}(x):=\tfrac{c_{3}x^{\rho}}{1+h^{\alpha}x^{2\kappa\alpha}},$ (3.9)
where $\alpha$ is a positive constant satisfying $\alpha\geq\tfrac{1}{2}$. Here the choice of $f_{h}$ and $g_{h}$ are nothing but a kind of taming strategy, which was firstly proposed in [10] and later extensively studied by many authors, see, e.g., [19, 3, 16, 17, 12, 13]. To validate Assumption 3.2, noticing that $2\kappa\alpha\geq\kappa>\rho$, simple calculations infer that for any $x\in(0,+\infty)$
$|f_{h}(x)|^{2\alpha}\leq\tfrac{c_{2}^{2\alpha}x^{2\kappa\alpha}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{2}^{2\alpha}h^{-\alpha},\quad|g_{h}(x)|^{\frac{2\kappa\alpha}{\rho}}\leq\tfrac{c_{3}^{\frac{2\kappa\alpha}{\rho}}x^{2\kappa\alpha}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{3}^{\frac{2\kappa\alpha}{\rho}}h^{-\alpha}.$ (3.10)
Thus we have
[…]
which confirms Assumption 3.2. To check Assumption 3.3, we consider two possibilities: $\kappa+1>2\rho$ and $\kappa+1=2\rho$. In the former case, one can easily observe that for any $\gamma>0$ there exists a constant $L>0$ such that
$\langle x,f(x)\rangle+\gamma|g(x)|^{2}$ $=-c_{2}x^{\kappa+1}+\gamma c_{3}^{2}x^{2\rho}\leq L,$ (3.12)
and thus
$\langle x,f_{h}(x)\rangle+\gamma|g_{h}(x)|^{2}=\tfrac{-c_{2}x^{\kappa+1}+\gamma c_{3}^{2}x^{2\rho}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq\tfrac{L}{1+h^{\alpha}x^{2\kappa\alpha}}\leq L.$ (3.13)
For the latter case $\kappa+1=2\rho$, one needs $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ to ensure the above two inequalities hold, which confirm Assumption 3.3. Finally, it can be deduced that
$\big{|}f(x)-f_{h}(x)\big{|}$ $=\tfrac{c_{2}h^{\alpha}x^{(2\alpha+1)\kappa}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{2}h^{\alpha}x^{(2\alpha+1)\kappa},$ (3.14)
$\big{|}g(x)-g_{h}(x)\big{|}$ $=\tfrac{c_{3}h^{\alpha}x^{(2\alpha+1)\rho}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{3}h^{\alpha}x^{(2\alpha+1)\rho}.$ (3.15)
Let $m_{1}=2(2\alpha+1)\kappa$ and $m_{2}=2(2\alpha+1)\rho$. Since $\alpha\geq\tfrac{1}{2}$, (3.8) holds true on an additional condition $12(2\alpha+1)\kappa\leq 2\gamma+$ 1.
In what follows we establish the integrability of the numerical approximations.
Lemma 3.6.
Let Assumptions 3.2, 3.3 hold. Then for any $12\leq p\leq 2\gamma+$ 1 and for any step-size $h>0$ it holds that
$\sup_{0\leq n\leq N}\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}\leq C(1+|x_{0}|^{p})<\infty.$ (3.16)
Proof: To begin with, the proposed scheme (3.1) can be rewritten as
$Z_{n+1}=Y_{n}+\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}h+g_{h}(Y_{n})\Delta W_{n},\quad n=0,1,...,N-1,$ (3.17)
where we denote
$Z_{n+1}:=Y_{n+1}-c_{-1}hY_{n+1}^{-1}.$ (3.18)
Next we introduce a time-continuous version of $Z_{n+1}$ in each time interval $[t_{n},t_{n+1}]$ as follows:
$\widetilde{Z}_{t}^{(n)}:=Y_{n}+\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}(t-t_{n})+g_{h}(Y_{n})(W_{t}-W_{t_{n}}),\quad t\in[t_{n},t_{n+1}].$ (3.19)
Note that $\widetilde{Z}_{t_{n}}^{(n)}=Y_{n}\neq Z_{n}=\widetilde{Z}_{t_{n}}^{(n-1)}$. Owing to (3.6), a rough moment bound can be derived for $\widetilde{Z}_{t}^{(n)}$. Indeed, by the Jensen inequality, for any $p\geq 1$ we have
$\sup_{t_{n}\leq t\leq t_{n+1}}\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}\widetilde{Z}_{t}^{(n)}\big{|}^{2}\Big{)}^{p}\bigg{]}\leq\hat{C}\mathbb{E}\Big{[}\big{|}Y_{n}\big{|}^{2p}\Big{]}+\tilde{C}_{\varepsilon},$ (3.20)
where $\varepsilon\geq 2c_{-1}h>0$ is a positive constant specified later and $\hat{C}$$,\tilde{C}_{\varepsilon}$ are positive constants independent of $h$. Thus
$\mathbb{E}\bigg{[}\Big{(}\varepsilon+|Y_{n+1}|^{2}+c_{-1}^{2}h^{2}|Y_{n+1}|^{-2}-2c_{-1}h\Big{)}^{p}\bigg{]}=\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}Y_{n+1}-c_{-1}hY_{n+1}^{-1}\big{|}^{2}\Big{)}^{p}\bigg{]}\leq\hat{C}\mathbb{E}\Big{[}\big{|}Y_{n}\big{|}^{2p}\Big{]}+\tilde{C}_{\varepsilon}.$ (3.21)
By noting $\varepsilon\geq 2c_{-1}h$, it holds that
$\mathbb{E}\big{[}\big{|}Y_{n+1}\big{|}^{2p}\big{]}\leq\hat{C}\mathbb{E}\big{[}\big{|}Y_{n}\big{|}^{2p}\big{]}+\tilde{C}_{\varepsilon}.$ (3.22)
By iteration we get
$\sup_{0\leq n\leq N}\mathbb{E}\big{[}\big{|}Y_{n}\big{|}^{2p}\big{]}<\infty,$ (3.23)
which, together with (3.20), infers that
$\sup_{0\leq n\leq N-1}\sup_{t_{n}\leq t\leq t_{n+1}}\mathbb{E}\big{[}\big{(}\varepsilon+\big{|}\widetilde{Z}_{t}^{(n)}\big{|}^{2}\big{)}^{p}\big{]}<\infty.$ (3.24)
We emphasize that the above rough moment bound, depending on $N$, is not proved to be uniformly bounded with respect to $N$. Nevertheless, it suffices to allow the use of the Gronwall lemma later to arrive at uniform moment bounds. To do this, we let $\varepsilon\geq 1+2c_{-1}h$ in the following and denote
$V_{\varepsilon,q}(x):=\big{(}\varepsilon+|x|^{2}\big{)}^{q},\quad q,x\in\mathbb{R}.$ (3.25)
It can be easily checked that for any $q,q^{\prime}\in\mathbb{R}$,
$V_{\varepsilon,q}(x)\cdot V_{\varepsilon,q^{\prime}}(x)=V_{\varepsilon,q+q^{\prime}}(x),\quad\big{(}V_{\varepsilon,q}(x)\big{)}^{q^{\prime}}=V_{\varepsilon,q\cdot q^{\prime}}(x).$ (3.26)
Since $\varepsilon\geq 1+2c_{-1}h>1$, one can directly check that
$V_{\varepsilon,q}\leq V_{\varepsilon,q^{\prime}},\quad q\leq q^{\prime}.$ (3.27)
Also, elementary calculations, for $q\neq 0$, show
$V_{\varepsilon,q}^{\prime}(x)$ $=2qx\big{(}\varepsilon+|x|^{2}\big{)}^{q-1}=2qxV_{\varepsilon,q-1}(x),$ (3.28)
$V_{\varepsilon,q}^{\prime\prime}(x)$ $=2q\big{(}\varepsilon+|x|^{2}\big{)}^{q-2}\big{[}\varepsilon+(2q-1)|x|^{2}\big{]}$ $\leq 2q\cdot\max{\{1,2q-1\}}\cdot V_{\varepsilon,q-1}(x).$
and that
$x\cdot V_{\varepsilon,q}(x)\leq\big{(}\varepsilon+|x|^{2}\big{)}^{\frac{1}{2}}\cdot V_{\varepsilon,q}(x)=V_{\varepsilon,\frac{1}{2}}(x)\cdot V_{\varepsilon,q}(x)=V_{\varepsilon,q+\frac{1}{2}}(x).$ (3.29)
Equipped with these properties, we can start the estimation of $\widetilde{Z}_{t}^{(n)}$. The Itô formula together with (3.28) and (3.29) yields
$V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})$ $=V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{t}\bigg{[}V_{\varepsilon,q}^{\prime}(\widetilde{Z}_{s}^{(n)})\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}+\tfrac{1}{2}V_{\varepsilon,q}^{\prime\prime}(\widetilde{Z}_{s}^{(n)})|g_{h}(Y_{n})|^{2}\bigg{]}\,\text{d}s$ (3.30)
$\quad+\int_{t_{n}}^{t}V_{\varepsilon,q}^{\prime}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}$
$\leq V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{t}\bigg{[}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}$
$\quad+q(2q-1)V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})|g_{h}(Y_{n})|^{2}\bigg{]}\,\text{d}s+\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}$
$\leq V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{t}2qV_{\varepsilon,q-\frac{1}{2}}(\widetilde{Z}_{s}^{(n)})\cdot\big{|}-c_{0}+c_{1}Y_{n}\big{|}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\Big{[}\big{\langle}Y_{n},f_{h}(Y_{n})\big{\rangle}+\tfrac{2q-1}{2}|g_{h}(Y_{n})|^{2}\Big{]}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\big{(}\widetilde{Z}_{s}^{(n)}-Y_{n}\big{)}f_{h}(Y_{n})\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}$
$=V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{t}2qV_{\varepsilon,q-\frac{1}{2}}(\widetilde{Z}_{s}^{(n)})\cdot\big{|}-c_{0}+c_{1}Y_{n}\big{|}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\Big{[}\big{\langle}Y_{n},f_{h}(Y_{n})\big{\rangle}+\tfrac{2q-1}{2}|g_{h}(Y_{n})|^{2}\Big{]}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s},$
where we also used (3.19) in the last step. In view of (3.7), for any $q\leq\gamma+\tfrac{1}{2}$, one deduces that
$\big{\langle}Y_{n},\,f_{h}(Y_{n})\big{\rangle}+\tfrac{2q-1}{2}|g_{h}(Y_{n})|^{2}\leq L,$ (3.31)
which, together with (3.6), further implies that
$f_{h}(Y_{n})\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}$ $\leq-c_{0}f_{h}(Y_{n})+c_{1}L+|f_{h}(Y_{n})|^{2}$ (3.32)
$\leq C(1+|f_{h}(Y_{n})|^{2})$
$\leq C(1+h^{-1}).$
Taking these estimates into account, one can derive from (3.27) and (3.30) that
$V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})$ $\leq V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{t}2qc_{0}V_{\varepsilon,q}(\widetilde{Z}_{s}^{(n)})\,\text{d}s+\int_{t_{n}}^{t}2qc_{1}V_{\varepsilon,q-\tfrac{1}{2}}(\widetilde{Z}_{s}^{(n)})\cdot Y_{n}\,\text{d}s$ (3.33)
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})(L+C)\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}.$
Applying the Young inequality twice leads to
$V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})$ $\leq V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{t}2qc_{0}V_{\varepsilon,q}(\widetilde{Z}_{s}^{(n)})\,\text{d}s+\int_{t_{n}}^{t}c_{1}(2q-1)V_{\varepsilon,q}(\widetilde{Z}_{s}^{(n)})\,\text{d}s+\int_{t_{n}}^{t}c_{1}|Y_{n}|^{2q}\,\text{d}s$ (3.34)
$\quad+\int_{t_{n}}^{t}(2q-2)V_{\varepsilon,q}(\widetilde{Z}_{s}^{(n)})\,\text{d}s+\int_{t_{n}}^{t}2(L+C)^{q}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s$
$\quad+\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}.$
Observing that
$\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}\Bigg{]}=0,$ (3.35)
taking expectations on both sides of (3.34) reveals that
$\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})\Big{]}$ $\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q(c_{0}+c_{1}+1)-c_{1}-2)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,q}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch$ (3.36)
$\quad+\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}.$
The case $q=1$ directly results in
$\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t}^{(n)})\Big{]}$ $\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+c_{1}\cdot\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,1}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch.$ (3.37)
Now we focus on the case $q>1$. Applying the Itô formula again on $V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})$ we get
$V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})$ $=V_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{s}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})\cdot\big{(}-c_{0}+c_{1}Y_{n}\big{)}\,\text{d}r$ (3.38)
$\quad+\int_{t_{n}}^{s}\bigg{[}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})f_{h}(Y_{n})+\tfrac{1}{2}V_{\varepsilon,q-1}^{\prime\prime}(\widetilde{Z}_{r}^{(n)})|g_{h}(Y_{n})|^{2}\bigg{]}\,\text{d}r$
$\quad+\int_{t_{n}}^{s}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})g_{h}(Y_{n})\,\text{d}W_{r}.$
Plugging (3.38) into (3.36) and noting that
$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}=0,$ (3.39)
we have
$\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})\Big{]}$ (3.40)
$\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q(c_{0}+c_{1}+1)-c_{1}-2)\cdot\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,q}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch$
$\quad+\underbrace{\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2q\int_{t_{n}}^{s}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})\cdot\big{(}-c_{0}+c_{1}Y_{n}\big{)}\,\text{d}r\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}}_{=:D_{1}}$
$\quad+\underbrace{\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2q\int_{t_{n}}^{s}\bigg{[}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})f_{h}(Y_{n})+\tfrac{1}{2}V_{\varepsilon,q-1}^{\prime\prime}(\widetilde{Z}_{r}^{(n)})|g_{h}(Y_{n})|^{2}\bigg{]}\,\text{d}r\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}}_{=:D_{2}}$
$\quad+\underbrace{\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2q\int_{t_{n}}^{s}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})g_{h}(Y_{n})\,\text{d}W_{r}\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}.}_{=:D_{3}}$
Using the Young inequality twice as well as the Jensen inequality on $D_{1}$ we get
$D_{1}$ $=4q(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\ \int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\cdot\big{(}-c_{0}+c_{1}Y_{n}\big{)}\,\text{d}r\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}$ (3.41)
$\leq 2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\bigg{(}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}$
$\quad+(2q-1)\Big{|}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\cdot\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{\frac{2q}{2q-1}}\bigg{)}\,\text{d}s\bigg{]}$
$\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\bigg{(}(2q-2)^{2}\Big{|}h^{-1}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{q}{q-1}}$
$\quad+(2q-2)\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\bigg{)}\,\text{d}s\bigg{]}$
$\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{q+1},$
where we used (3.6), (3.27), (3.29), the Hölder inequality and the moment inequality in the last step. Since the estimation of $D_{3}$ is similar and easy, we present it before treating $D_{2}$. By the generalized Itô isometry [8, Chapter 4, B], the Young inequality as well as (3.6), (3.27), (3.29), the Hölder inequality and the moment inequality we deduce that
$D_{3}$ $=4q(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})f_{h}(Y_{n})|g_{h}(Y_{n})|^{2}\,\text{d}r\,\text{d}s\bigg{]}$ (3.42)
$\leq 4(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\bigg{(}(q-1)\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{q}{q-1}}+\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})|g_{h}(Y_{n})|^{2}\,\text{d}r\Big{|}^{q}\bigg{)}\,\text{d}s\bigg{]}$
$\leq(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{\frac{q}{2}+1}.$
As for $D_{2}$, noting $\max{\{1,2q-3\}}<2q-1$ and using (3.28), one can easily show that for $q>1$,
$D_{2}$ $\leq 2q\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{|}\int_{t_{n}}^{s}\Big{(}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})f_{h}(Y_{n})+\tfrac{1}{2}V_{\varepsilon,q-1}^{\prime\prime}(\widetilde{Z}_{r}^{(n)})|g_{h}(Y_{n})|^{2}\Big{)}\,\text{d}r\bigg{|}\cdot\bigg{|}\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\bigg{|}\,\text{d}s\Bigg{]}$ (3.43)
$\leq 2q\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{|}\int_{t_{n}}^{s}\Big{(}2(q-1)\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})f_{h}(Y_{n})$
$\qquad+(q-1)\max{\{1,2q-3\}}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})|g_{h}(Y_{n})|^{2}\Big{)}\,\text{d}r\bigg{|}\cdot\bigg{|}\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\bigg{|}\,\text{d}s\Bigg{]}$
$<4q(q-1)\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{|}\int_{t_{n}}^{s}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\Big{[}\big{\langle}Y_{n},f_{h}(Y_{n})\big{\rangle}+\tfrac{2q-1}{2}|g_{h}(Y_{n})|^{2}\Big{]}\,\text{d}r\bigg{|}$
×$\quad\times\bigg{|}\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\bigg{|}\,\text{d}s\Bigg{]}$
$\quad+4q(q-1)\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{|}\int_{t_{n}}^{s}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\big{(}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{)}f_{h}(Y_{n})\,\text{d}r\bigg{|}\cdot\bigg{|}\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\bigg{|}\,\text{d}s\Bigg{]}$
$=:D_{21}+D_{22}.$
Employing the Young inequality, (3.27), (3.31) on $D_{21}$ as well as the Hölder inequality and the moment inequality we get
$D_{21}$ $\leq 4Lq(q-1)\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{(}\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}\cdot\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}\bigg{)}\,\text{d}s\Bigg{]}$ (3.44)
$\leq 4L(q-1)\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{(}(q-1)\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{q}{q-1}}+\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{q}\bigg{)}\,\text{d}s\Bigg{]}$
$\leq L(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{\frac{q}{2}+1}.$
With regard to $D_{22}$, using the Young inequality twice again as well as the Hölder inequality, the moment inequality, (3.6), (3.19) and (3.27) on $D_{22}$ we deduce that
$D_{22}$ $\leq 2(q-1)\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{(}(2q-1)\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})(\widetilde{Z}_{r}^{(n)}-Y_{n})\,\text{d}r\Big{|}^{\frac{2q}{2q-1}}$ (3.45)
$\quad+\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})^{2}g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\bigg{)}\,\text{d}s\Bigg{]}$
$\leq\tfrac{2q-2}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}(2q-1)\mathbb{E}\bigg{[}\big{|}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{2q}{2q-1}}\cdot\big{|}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{|}^{\frac{2q}{2q-1}}\bigg{]}\,\text{d}r\text{d}s+Ch$
$\leq\tfrac{2q-2}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\bigg{[}\big{|}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{2q-2}{2q-1}}\cdot\big{|}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{|}^{\frac{2q}{2q-1}}\bigg{]}\,\text{d}r\text{d}s+Ch$
$\leq\tfrac{(2q-2)^{2}}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\,\text{d}r\text{d}s+Ch$
$\quad+\tfrac{(2q-1)(2q-2)}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}+\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}\,\text{d}r\text{d}s$
$\leq(2q-2)^{2}h\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+C\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch.$
Combining all the above estimations and taking supremum reveal that
$\sup_{t_{n}\leq r\leq t}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}$ $\leq(1+Ch)\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+C\cdot\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\,\text{d}s+Ch.$ (3.46)
[…]
$\sup_{t_{n}\leq t\leq t_{n+1}}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})\Big{]}\leq\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}\text{e}^{Ch}+Ch\text{e}^{Ch},$ (3.47)
which directly infers that
$\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n+1}}^{(n)})\Big{]}=\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n+1}-c_{-1}hY_{n+1}^{-1}|^{2}\big{)}^{q}\Big{]}\leq\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}\Big{]}\text{e}^{Ch}+Ch\text{e}^{Ch}.$ (3.48)
From now on, we set $\varepsilon=1+c_{-1}+2c_{-1}h$. On the one hand, one knows that
$\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n+1}-c_{-1}hY_{n+1}^{-1}|^{2}\big{)}^{q}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}1+c_{-1}+2c_{-1}h+|Y_{n+1}|^{2}-2c_{-1}h+|Y_{n+1}|^{-2}\big{)}^{q}\Big{]}$ (3.49)
$\geq\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n+1}|^{2}\big{)}^{q}\Big{]}.$
On the other hand,
$\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}1+c_{-1}+2c_{-1}h+|Y_{n}|^{2}\big{)}^{q}\Big{]}$ (3.50)
$\leq(1+2h)^{q}\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n}|^{2}\big{)}^{q}\Big{]}$
$\leq\text{e}^{2hq}\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n}|^{2}\big{)}^{q}\Big{]}.$
Thus we conclude that
$\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n+1}|^{2}\big{)}^{q}\Big{]}\leq\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n}|^{2}\big{)}^{q}\Big{]}\text{e}^{Ch}+Che^{Ch}.$ (3.51)
By iteration we arrive at
$\sup_{1\leq n\leq N}\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n}|^{2})^{q}\Big{]}$ $\leq\sup_{1\leq n\leq N}\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{0}|^{2}\big{)}^{q}\Big{]}\text{e}^{Cnh}+C<\infty,$ (3.52)
as required. ∎
Equipped with the assumptions on the modifications and the integrability of numerical solutions, we can derive an important result on the bounds of modification errors.
Corollary 3.7.
Let Assumptions 3.2, 3.3, 3.4 hold. Then it holds that
$\mathbb{E}\Big{[}\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\Big{]}\leq Ch,\quad\mathbb{E}\Big{[}\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\Big{]}\leq Ch.$ (3.53)
The above properties of the numerical approximations play an important role in the following error analysis.
4 Mean-square convergence rate of the proposed scheme
In this section we present the analysis of the mean-square convergence rate for the proposed scheme on the mesh grid. Two cases are considered including the non-critical case $\kappa+1>2\rho$ and the general critical case $\kappa+1=2\rho$. For both cases, we successfully identify the desired mean-square convergence rate of order one-half for the considered scheme. The main convergence result is formulated as follows.
Theorem 4.1.
Let $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{n}\}_{0\leq n\leq N}$ be defined by (2.2) and (3.1), respectively. Let Assumptions 3.2, 3.3 and 3.4 hold, with additionally one of the following conditions standing:
  1. (1)
    $\kappa+1>2\rho$;
  2. (2)
    $\kappa+1=2\rho$, $m_{1},m_{2}\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$, $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$.
Then there exists a constant $C>0$ independent of $h>0$ such that
$\sup_{0\leq n\leq N}\mathbb{E}\big{[}\big{|}X_{t_{n}}-Y_{n}\big{|}^{2}\big{]}\leq Ch.$ (4.1)
Proof: To carry out the error analysis, we first obtain the error recurrence formula. By recalling definitions of $X_{t}$, $Y_{n}$, $f$, $g$ in (2.2), (3.1), (3.2) and denoting
$e_{n}:=X_{t_{n}}-Y_{n},\quad n=0,1,...,N$
for brevity, one can straightforwardly infer
$e_{n+1}-c_{-1}\big{(}X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{)}h$ $=(1+c_{1}h)e_{n}+\big{(}f(X_{t_{n}})-f(Y_{n})\big{)}h+\big{(}g(X_{t_{n}})-g(Y_{n})\big{)}\Delta W_{n}+R_{n+1},$ (4.2)
where
$R_{n+1}$ $:=\int_{t_{n}}^{t_{n+1}}\Big{(}c_{-1}X_{s}^{-1}+c_{1}X_{s}+f(X_{s})\Big{)}\,\text{d}s+\int_{t_{n}}^{t_{n+1}}g(X_{s})\,\text{d}W_{s}-c_{1}X_{t_{n}}h-c_{-1}X_{t_{n+1}}^{-1}h$ (4.3)
$\quad-\big{(}f(X_{t_{n}})-f(Y_{n})\big{)}h-\big{(}g(X_{t_{n}})-g(Y_{n})\big{)}\Delta W_{n}-f_{h}(Y_{n})h-g_{h}(Y_{n})\Delta W_{n}$
$=\underbrace{\int_{t_{n}}^{t_{n+1}}\Big{[}c_{-1}\big{(}X_{s}^{-1}-X_{t_{n+1}}^{-1}\big{)}+c_{1}(X_{s}-X_{t_{n}})+\big{(}f(X_{s})-f(X_{t_{n}})\big{)}+(f(Y_{n})-f_{h}(Y_{n}))\Big{]}\,\text{d}s}_{=:R_{n+1}^{(1)}}$
$\quad+\underbrace{\int_{t_{n}}^{t_{n+1}}\Big{[}\big{(}g(X_{s})-g(X_{t_{n}})\big{)}+\big{(}g(Y_{n})-g_{h}(Y_{n})\big{)}\Big{]}\,\text{d}W_{s}}_{=:R_{n+1}^{(2)}}.$
Squaring both sides of (4.2) and using the Young inequality four times yield
$|e_{n+1}|^{2}-2c_{-1}h\big{\langle}e_{n+1},X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{\rangle}+c_{-1}^{2}h^{2}\big{|}X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{|}^{2}$ (4.4)
$=(1+c_{1}h)^{2}|e_{n}|^{2}+|f(X_{t_{n}})-f(Y_{n})|^{2}h^{2}+|g(X_{t_{n}})-g(Y_{n})|^{2}|\Delta W_{n}|^{2}+|R_{n+1}|^{2}$
$\quad+2h(1+c_{1}h)\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle+2(1+c_{1}h)\langle e_{n},(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\rangle$
$\quad+2(1+c_{1}h)\langle e_{n},R_{n+1}\rangle+2h\big{\langle}f(X_{t_{n}})-f(Y_{n}),(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\big{\rangle}$
$\quad+2h\big{\langle}f(X_{t_{n}})-f(Y_{n}),R_{n+1}\big{\rangle}+2\big{\langle}(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n},R_{n+1}\big{\rangle}$
$\leq(1+c_{1}h)^{2}|e_{n}|^{2}+|f(X_{t_{n}})-f(Y_{n})|^{2}h^{2}+|g(X_{t_{n}})-g(Y_{n})|^{2}|\Delta W_{n}|^{2}+|R_{n+1}|^{2}$
$\quad+2h\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle+c_{1}h|e_{n}|^{2}+c_{1}h^{3}|f(X_{t_{n}})-f(Y_{n})|^{2}$
$\quad+2(1+c_{1}h)\langle e_{n},(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\rangle+2(1+c_{1}h)\langle e_{n},R_{n+1}^{(2)}\rangle+h(1+c_{1}h)^{2}|e_{n}|^{2}+\tfrac{1}{h}|R_{n+1}^{(1)}|^{2}$
$\quad+2h\big{\langle}f(X_{t_{n}})-f(Y_{n}),(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\big{\rangle}+h^{2}|f(X_{t_{n}})-f(Y_{n})|^{2}+|R_{n+1}|^{2}$
$\quad+\delta|g(X_{t_{n}})-g(Y_{n})|^{2}|\Delta W_{n}|^{2}+\tfrac{1}{\delta}|R_{n+1}|^{2},$
where $\delta>0$ is some positive constant. By noting that
$\big{\langle}e_{n+1},X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{\rangle}=\int_{0}^{1}-\Big{(}Y_{n+1}+\theta(X_{t_{n+1}}-Y_{n+1})\Big{)}^{-2}\text{d}\theta\cdot(X_{t_{n+1}}-Y_{n+1})^{2}\leq 0$ (4.5)
and that
$\mathbb{E}\big{[}\langle e_{n},(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\rangle\big{]}=0,\,\mathbb{E}\big{[}\langle e_{n},R_{n+1}^{(2)}\rangle\big{]}$ $=0,$
$\mathbb{E}\big{[}\big{\langle}f(X_{t_{n}})-f(Y_{n}),(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\big{\rangle}\big{]}$ $=0,$
we thus take expectations on both sides of (4.4) and do some rearrangements to arrive at
$\mathbb{E}\big{[}|e_{n+1}|^{2}\big{]}$ $\leq(1+(1+3c_{1})h+(c_{1}^{2}+2c_{1})h^{2}+c_{1}^{2}h^{3})\mathbb{E}\big{[}|e_{n}|^{2}\big{]}+(2+c_{1}h)h^{2}\mathbb{E}\big{[}|f(X_{t_{n}})-f(Y_{n})|^{2}\big{]}$ (4.6)
$\quad+(1+\delta)h\mathbb{E}\big{[}|g(X_{t_{n}})-g(Y_{n})|^{2}\big{]}+2h\mathbb{E}\big{[}\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle\big{]}$
$\quad+(2+\tfrac{1}{\delta})\mathbb{E}\big{[}|R_{n+1}|^{2}\big{]}+\tfrac{1}{h}\mathbb{E}\big{[}|R_{n+1}^{(1)}|^{2}\big{]}.$
It is not difficult to check that
$(1+\delta)\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}+2\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle$ (4.7)
$=(1+\delta)c_{3}^{2}(X_{t_{n}}^{\rho}-Y_{n}^{\rho})^{2}-2c_{2}(X_{t_{n}}-Y_{n})(X_{t_{n}}^{\kappa}-Y_{n}^{\kappa})$
$=(1+\delta)c_{3}^{2}(X_{t_{n}}-Y_{n})^{2}\bigg{(}\int_{0}^{1}\rho(Y_{n}+\theta(X_{t_{n}}-Y_{n}))^{\rho-1}\,\text{d}\theta\bigg{)}^{2}$
$\quad-2c_{2}(X_{t_{n}}-Y_{n})^{2}\int_{0}^{1}\kappa(Y_{n}+\theta(X_{t_{n}}-Y_{n}))^{\kappa-1}\,\text{d}\theta$
$\leq(X_{t_{n}}-Y_{n})^{2}\int_{0}^{1}\big{[}(1+\delta)c_{3}^{2}\rho^{2}(\theta X_{t_{n}}+(1-\theta)Y_{n})^{2\rho-2}-2c_{2}\kappa(\theta X_{t_{n}}+(1-\theta)Y_{n})^{\kappa-1}\big{]}\,\text{d}\theta.$
In view of this, we claim that
$(1+\delta)h\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}+2h\mathbb{E}\big{[}\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle\big{]}\leq\tilde{L}(X_{t_{n}}-Y_{n})^{2}.$ (4.8)
Indeed, for the case $\kappa+1>2\rho$, the assertion is obvious by noting $\kappa-1>2\rho-2$ and $02c_{2}\kappa>$ 0. For the critical case $\kappa+1=2\rho$, the assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ implies $\tfrac{c_{2}}{c_{3}^{2}}>\kappa/2$ and thus $\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-1>0$. As a result, one can choose $10<\delta\leq\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-$ 1 such that $(1+\delta)c_{3}^{2}\rho^{2}-2c_{2}\kappa\leq 0$, which ensures that (4.8) still remains valid. Taking (4.8) into account, one can deduce from (4.6) that
$\mathbb{E}\big{[}|e_{n+1}|^{2}\big{]}$ $\leq(1+(1+3c_{1})h+(c_{1}^{2}+2c_{1})h^{2}+c_{1}^{2}h^{3})\mathbb{E}\big{[}|e_{n}|^{2}\big{]}+(2+c_{1}h)h^{2}\mathbb{E}\big{[}|f(X_{t_{n}})-f(Y_{n})|^{2}\big{]}$ (4.9)
$\quad+(2+\tfrac{1}{\delta})\mathbb{E}\big{[}|R_{n+1}|^{2}\big{]}+\tfrac{1}{h}\mathbb{E}\big{[}|R_{n+1}^{(1)}|^{2}\big{]}.$
Before proceeding further, two cases $\kappa+1>2\rho$ and $\kappa+1=2\rho$ are further considered separately as follows.
  1. Case 1:
    $\kappa+1>2\rho$.
    Due to Lemma 2.2 and Lemma 3.6, one can infer that
    $\mathbb{E}\big{[}\big{|}f(X_{t_{n}})-f(Y_{n})\big{|}^{2}\big{]}\leq C\quad\text{and}\quad\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}\leq C.$ (4.10)
    Corollary 3.7 implies that
    $\mathbb{E}\big{[}\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch,\quad\mathbb{E}\big{[}\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch.$ (4.11)
    Equipped with (4.11) as well as Lemmas 2.4, 2.6 one can directly obtain
    $\mathbb{E}\big{[}\big{|}R_{n+1}^{(1)}\big{|}^{2}\big{]}\leq Ch^{3},\quad\mathbb{E}\big{[}\big{|}R_{n+1}^{(2)}\big{|}^{2}\big{]}\leq Ch^{2}.$ (4.12)
  2. Case 2:
    $\kappa+1=2\rho.$
    The difference for the critical case comes from the extra restriction that $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$, as required by Assumption 3.3, which together with Lemmas 2.3, 3.6 ensures the moment boundedness of $\mathbb{E}\big{[}|X_{t}|^{p}\big{]}$ and $\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}$ for $p\leq 2\gamma+1\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$. Therefore, on the condition $\tfrac{c_{2}}{c_{3}^{2}}>\kappa-\tfrac{1}{2}$, one can derive $2\kappa$-th moment bounds for both analytic and numerical solutions
    $\mathbb{E}\big{[}\big{|}f(X_{t_{n}})-f(Y_{n})\big{|}^{2}\big{]}\leq C\quad\text{and}\quad\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}\leq C.$ (4.13)
    Recalling (3.8) one knows that
    $\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\leq Ch|Y_{n}|^{m_{1}},\quad\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\leq Ch|Y_{n}|^{m_{2}},$ (4.14)
    where $m_{1},m_{2}\leq 2\gamma+1\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$. Thus for $m_{1},m_{2}\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$,
    $\mathbb{E}\big{[}\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch,\quad\mathbb{E}\big{[}\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch,$ (4.15)
    by Corollary 3.7. This together with Lemma 2.5 and 2.6 helps us to deduce that, on the more strict assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$,
    $\mathbb{E}\big{[}\big{|}R_{n+1}^{(1)}\big{|}^{2}\big{]}\leq Ch^{3},\quad\mathbb{E}\big{[}\big{|}R_{n+1}^{(2)}\big{|}^{2}\big{]}\leq Ch^{2}.$ (4.16)
Based on the above discussion, one can derive from (4.9) that
$\mathbb{E}\big{[}|e_{n+1}|^{2}\big{]}$ $\leq(1+Ch)\mathbb{E}\big{[}|e_{n}|^{2}\big{]}+Ch^{2},$ (4.17)
which, by iteration gives
$\sup_{0\leq n\leq N}\mathbb{E}\big{[}|e_{n}|^{2}\big{]}\leq Ch,$ (4.18)
as required. ∎
Theorem 4.1 can be applied to derive convergence rates for any scheme satisfying Assumptions 3.2, 3.3 and 3.4. The following corollary shows an application on the schemes in Example 3.5.
Corollary 4.2 (Convergence rates of semi-implicit tamed Euler methods).
Let $\{X_{t}\}_{0\leq t\leq T}$ and $\{Y_{n}\}_{0\leq n\leq N}$ be defined as (2.2) and (3.1), respectively. Let $f_{h}$ and $g_{h}$ be defined as in Example 3.5. Suppose that Assumptions 3.2, 3.3 and 3.4 hold, with additionally one of the following conditions standing:
  1. (1)
    $\kappa+1>2\rho$.
  2. (2)
    $\kappa+1=2\rho$, $\tfrac{c_{2}}{c_{3}^{2}}\geq(2\alpha+1)\kappa-\tfrac{1}{2}$.
Then there exists a constant $C>0$ independent of the step-size $h$ such that
$\sup_{0\leq n\leq N}\mathbb{E}\big{[}\big{|}X_{t_{n}}-Y_{n}\big{|}^{2}\big{]}\leq Ch.$ (4.19)
We would like to mention that, for the general critical case $\kappa+1=2\rho$ one can observe a slightly more strict restriction on the model parameters than that for the implicit methods (STMs) in [20], where $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ is required. This is caused by the need of $m_{1},m_{2}\leq 2\gamma+1$ in order to confirm Assumption 3.4. More precisely, as indicated by Example 3.5, for the general critical case, $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ as well as $12(2\alpha+1)\rho=m_{2}<m_{1}=2(2\alpha+1)\kappa\leq 2\gamma+$ 1 are required, leading to $(2\alpha+1)\kappa\leq\tfrac{c_{2}}{c_{3}^{2}}+\tfrac{1}{2}$ as mentioned.
5 Numerical experiments
In this section, numerical experiments are presented to illustrate the previous theoretical results. Let us consider the following Aït-Sahalia model
$\mathrm{d}X_{t}=$ $(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\mathrm{d}t+c_{3}X_{t}^{\rho}\mathrm{d}W_{t},\quad X_{0}=1>0.$ (5.1)
The semi-implicit tamed Euler-type method (TEM) in Example 3.5 with $\alpha=\tfrac{1}{2}$ is applied to approximate the the model:
$Y_{n+1}=Y_{n}+c_{-1}Y_{n+1}^{-1}h+\big{(}-c_{0}+c_{1}Y_{n}-\tfrac{c_{2}Y_{n}^{\kappa}}{1+\sqrt{h}Y_{n}^{\kappa}}\big{)}h+\tfrac{c_{3}Y_{n}^{\rho}}{1+\sqrt{h}Y_{n}^{\kappa}}\Delta W_{n}.$ (5.2)
The following two sets of parameters are taken as numerical tests.
Example 5.1.
$\kappa=5,\rho=1.5,c_{-1}=1.5,c_{0}=2,c_{1}=1,c_{2}=2,c_{3}=1$.
Example 5.2.
$\kappa=2,\rho=1.5,c_{-1}=1.5,c_{0}=2,c_{1}=1,c_{2}=4,c_{3}=1$.
Note that Example 5.1 satisfies the non-critical condition $\kappa+1>2\rho$, while Example 5.2 satisfies the general critical condition $\kappa+1=2\rho$. The strong convergence rates are tested with fixed total time $T=1$. The numerical approximations with the step-size $h_{exact}=2^{-14}$ is used to simulate the ”exact” solutions, while $h=2^{-i},i=4,5,...,9$ are used for numerical solutions over $10^{4}$ Brownian paths. The estimation error is calculated in terms of $e_{h}:=\bigg{(}\sup\limits_{0\leq n\leq N}\mathbb{E}\Big{[}|X_{t_{n}}-Y_{n}|^{2}\Big{]}\bigg{)}^{\tfrac{1}{2}}$. The stochastic theta method (STM) [20] with $\theta=1$, which coincides with the backward Euler-Maruyama method (BEM) [18], is also tested in comparison with TEM.
Fig. 1 plots the average sample error against various step-sizes on a log-log scale, with BEM in a black solid line and TEM in a red solid line. It can be easily observed the order $0.5$ strong convergence for both two schemes. By a least-squares fitting, the resulting convergence rate $q$ and the least squares residual are shown in Table 1, where one can also confirm the predicted convergence rate.
Moreover, the computational costs of BEM and the proposed TEM scheme over $10^{4}$ Brownian paths are shown in Table 2 and 50 out of the $10^{4}$ paths are shown in Fig. 2. It is easy to see that the TEM can preserve positivity and reduce computational costs sigficantly, compared with BEM, especially as the parameter $\kappa$ increases.
Figure 1: Strong convergence rate of the semi-implicit taming Euler method under Example 5.1 (left) and Example 5.2 (right).
Figure 2: 50 paths of numerical solutions for BEM and TEM
Table 1: A least squares fit for the convergence rate $q$
BEM TEM
Case 1 $q=0.5409$, $resid=0.0269$ $q=0.5173$, $resid=0.0128$
Case 2 $q=0.5558$, $resid=0.0698$ $q=0.5288$, $resid=0.0326$
Table 2: Time cost (seconds) for BEM and TEM over $10^{4}$ Brownian paths
BEM TEM
Example 5.1 99.937 0.193
Example 5.2 47.997 0.094
6 Conclusion
In this paper we propose and analyze an explicit time-stepping scheme for the strong approximations of the generalized Aït-Sahalia model. Based on a semi-implicit strategy and proper modifications, the numerical scheme is designed to be positivity preserving and easily implementable with cheap computational costs compared with existing positivity preserving schemes. The mean-square strong convergence rate of the scheme is carefully analyzed and proved to be order $0.5$. Numerical experiments are also provided to support the theoretical findings. As an ongoing project, we successfully construct higher order efficient positivity preserving time-stepping schemes for the Aït-Sahalia model [11] and try to cover a larger range of stochastic models [15].
Declarations
Ethical Approval
Not applicable
Availability of supporting data
The authors confirm that the data supporting the findings of this study are available within the article.
Competing interests
The authors declare no competing interests.
Funding
This work is supported by the National Natural Science Foundation of China (Grant Nos. 12071488, 12371417, 11971488), the Natural Science Foundation of Hunan Province (Grant No.2020JJ2040).
Authors’ contributions
All the authors contribute equally to this work.
Acknowledgments
Not applicable

### v2

@@
 Our work is motivated by the need to justify the multi-level Monte Carlo (MLMC) simulations for the underlying model, where the rate of mean-square convergence is required and the preservation of positivity is desirable particularly for large discretization time steps.
-To the best of our knowledge, this is the first paper to propose an unconditionally positivity preserving explicit scheme with order $1/2$ of mean-square convergence for the model.
 Numerical experiments are finally provided to confirm the theoretical findings.
-Keywords: Aït-Sahalia model; Explicit scheme; Unconditionally positivity preserving; Strong convergence rates.
+Keywords: Aït-Sahalia model; Explicit scheme; Unconditionally positivity preserving; Strong convergence rate.
 1 Introduction
 The pricing of financial assets has been an important issue in mathematical finance over the past few decades, where stochastic differential equations (SDEs) are prominent models frequently used to capture the behaviors of financial systems.
-Notable examples include the Black-Scholes model [2], the Merton model [14] and the Cox-Ingersoll-Ross (CIR) model [5] in the modern financial theory.
+Notable examples include the Black-Scholes model [2], the Merton model [16], and the Cox-Ingersoll-Ross (CIR) model [5] in the modern financial theory.
 The economist and mathematician Aït-Sahalia conducted an empirical study in [1] and investigated several continuous-time models for interest rates.
-He tested the models by comparing the parameter density implied by the parameter models, and found that all the existing univariate linear drift models fail to explain the euro dollar well.
+He tested the models by comparing the parameter density implied by the parameter models and found that all the existing univariate linear drift models fail to explain the euro dollar well.
 As a consequence of this work, a nonlinear stochastic interest model which captures well of the dynamics of the spot rate was proposed and named as the Aït-Sahalia model, which is now widely used in finance and economics, and described by the following Itô-type SDE:
 $\mathrm{d}X_{t}=\big{(}c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{2}\big{)}\,\text{d}t+c_{3}X_{t}^{\rho}\,\text{d}W_{t},\quad X_{0}=x_{0}>0.$ (1.1)
-Here $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\rho>1$ and $\{W_{t}\}_{t\in[0,T]}$ stands for a standard Brownian motion defined on a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$.
+Here, $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\rho>1$ and $\{W_{t}\}_{t\in[0,T]}$ stands for a standard Brownian motion defined on a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$.
 Aït-Sahalia [1] proved that the model (1.1) admits a unique global positive solution using the Feller explosion test.
@@
 [20] extended the model (1.1) to a more general version given by
-$\mathrm{d}X_{t}=$ $(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\text{d}t+c_{3}X_{t}^{\rho}\text{d}W_{t},\quad X_{0}=x_{0}>0,\>\kappa,\rho>1,$ (1.2)
+$\mathrm{d}X_{t}=$ $(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\text{d}t+c_{3}X_{t}^{\rho}\text{d}W_{t},\quad X_{0}=x_{0}>0,\>\kappa>1,$ (1.2)
 and showed that the generalized model admits a unique global positive solution with uniformly in time bounded $p$-th moments ($p\in\mathbb{R}$) in the non-critical case $\kappa+1>2\rho$.
@@
 The work [14] proposed an exponential tamed method based on the Lamperti transform and splitting, which is strongly convergent with order one, but not strictly positivity preserving.
-In the present article we aim to design novel explicit positivity preserving schemes for approximations of the Aït-Sahalia model (1.2) with mean-square convergence rate of order $\tfrac{1}{2}$.
+Finally, it is worthwhile to mention an interesting paper [10], where, by combining a Lamperti-type transformation with a semi-implicit approach, the authors proposed a positivity preserving scheme for the model in the particular critical case $\kappa=2,\rho=\tfrac{3}{2}$.
+For the particular critical case, the proposed scheme can be explicitly solved, by finding a positive root of a quadratic equation.
+Moreover, it was shown to have order one strong convergence (see [10] for more details).
+In the present article, we aim to design novel explicit positivity preserving schemes for approximations of the Aït-Sahalia model (1.2) with a mean-square convergence rate of order $\tfrac{1}{2}$.
 More precisely, on a uniform mesh with a uniform step-size $h=\tfrac{T}{N},\,N\in\mathbb{N}$ we propose the following time-stepping scheme:
 $Y_{n+1}=Y_{n}+c_{-1}Y_{n+1}^{-1}h+\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}h+g_{h}(Y_{n})\Delta W_{n},\quad Y_{0}=X_{0},$ (1.3)
-where $f_{h}$ and $g_{h}$ are some explicit modifications of $f(x):=-c_{2}x^{\kappa}$ and $g(x):=c_{3}x^{\rho}$ in the SDE model, respectively (see Section 3 for details).
+where $f_{h}$ and $g_{h}$ are some explicit modifications of $f(x):=-c_{2}x^{\kappa}$ and $g(x):=c_{3}x^{\rho}$ in the SDE model, respectively (see Sect.
+3 for details).
 Clearly, the implicitness only appears in the term $c_{-1}Y_{n+1}^{-1}h$, which is the key element to preserve the positivity of the original model (see Lemma 3.1).
@@
 Recall that the analysis of convergence rates in [22] for the implicit schemes does not rely on a priori moment estimates of numerical approximations.
-Nevertheless, due to the presence of the explicit treatment in the drift term, the analysis of convergence rates for the proposed semi-implicit schemes relies on finite moment bounds of numerical approximations:
+Nevertheless, due to the presence of the explicit treatment in the drift term, the analysis of convergence rates for the proposed semi-implicit schemes relies on a finite moment bound of numerical approximations:
 $\sup_{N\in\mathbb{N}}\sup_{0\leq n\leq N}\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}<\infty,\quad p\geq 2.$ (1.4)
-Deriving such $p$-th moment integrability of the numerical approximations is, however, not an easy task and some non-standard arguments are exploited (see the proof of Lemma 3.6), to overcome essential difficulties caused by a drift that blows up at the origin, highly nonlinear diffusion coefficients and a mixture of implicitness and explicitness in the drift part of the scheme.
+Deriving such $p$-th moment integrability of the numerical approximations is, however, not an easy task and some non-standard arguments are exploited (see the proof of Lemma 3.6) to overcome essential difficulties caused by a drift that blows up at the origin, highly nonlinear diffusion coefficients and a mixture of implicitness and explicitness in the drift part of the scheme.
 Equipped with the moment bound (1.4), one is able to identify the mean-square convergence rate of order $0.5$ for the proposed scheme (3.1) applied to the model (1.2) (see Theorem 4.1):
@@
 in both non-critical ($\kappa+1>2\rho$) and general critical ($\kappa+1=2\rho$) cases, where the constant $C$ is independent of the time step-size $h$.
-To the best of our knowledge, this is the first paper to devise an unconditionally positivity preserving explicit scheme with order $\tfrac{1}{2}$ of mean-square convergence for the general Aït-Sahalia model.
+To the best of our knowledge, this is the first paper to devise an unconditionally positivity preserving explicit scheme with order $\tfrac{1}{2}$ of mean-square convergence for the non-critical case of the Aït-Sahalia model (1.2).
 The rest of this paper is organized as follows.
 In the next section, basic notations and properties of the original model are presented.
-In Section 3 the newly proposed scheme is introduced and proved to be positivity preserving and moment bounded.
-The mean-square error bounds for the proposed scheme is carefully analyzed in Section 4, with order one-half of convergence revealed.
-Numerical experiments are shown in Section 5 to verify our theoretical results and a short conclusion is made in Section 6.
+In Sect.
+3 the newly proposed scheme is introduced and proved to be positivity preserving and moment bounded.
+The mean-square error bounds for the proposed scheme is carefully analyzed in Sect.
+4, with order one-half of convergence revealed.
+Numerical experiments are shown in Sect.
+5 to verify our theoretical results, and a short conclusion is made in Sect.
+6.
 2 The generalized Aït-Sahalia model
 […]
-$\|\xi\|_{\mathcal{L}^{p}(\Omega;\mathbb{R})}:=\big{(}\mathbb{E}\big{[}|\xi|^{p}\big{]}\big{)}^{1/p}<\infty.$ (2.1)
-In this work we focus on the generalized Aït-Sahalia model taking the form as follows:
+$\|\xi\|_{L^{p}(\Omega;\mathbb{R})}:=\big{(}\mathbb{E}\big{[}|\xi|^{p}\big{]}\big{)}^{1/p}<\infty.$ (2.1)
+In this work, we focus on the generalized Aït-Sahalia model taking the form as follows:
 $\text{d}X_{t}=(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\text{d}t+c_{3}X_{t}^{\rho}\text{d}W_{t},\quad X_{0}=x_{0},\quad t>0,$ (2.2)
@@
 For the particular case $\kappa=2$, the underlying model (2.2) reduces into (1.1), originally proposed by Aït-Sahalia [1].
-The well-posedness of the generalized model (2.2) has been established in the following theorem quoted from [18, Theorem 2.1].
+The well-posedness of the generalized model (2.2) has been established in the following theorem quoted from [20, Theorem 2.1].
 Theorem 2.1.
 Let $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$. Given any initial value $X_{0}>0$, there exists a unique, positive global solution $X_{t}$ to the SDE (2.2) on $t\geq 0$.
-Furthermore, we recall the moment integrability properties of the solution to (2.2) as follows (see [18, Lemma 2.1] and [20, Lemma 4.6] for details).
+Furthermore, we recall the moment integrability properties of the solution to (2.2) as follows (see [20, Lemma 2.1] and [22, Lemma 4.6] for details).
 Lemma 2.2.
 Let conditions in Theorem 2.1 be satisfied with $\kappa+1>2\rho$. For any $p\geq 2$, it holds
-[…]
-For the general critical case $\kappa+1=2\rho$, we have slightly different assertions, as stated in the following lemma (cf. [20, Lemma 4.6]).
+$\sup_{0\leq t<\infty}\mathbb{E}\big{[}|X_{t}|^{p}\big{]}<\infty\quad\text{and}\quad\sup_{0\leq t<\infty}\mathbb{E}\big{[}\big{|}\tfrac{1}{X_{t}}\big{|}^{p}\big{]}<\infty.$ (2.3)
+For the general critical case $\kappa+1=2\rho$, we have slightly different assertions, as stated in the following lemma (cf. [22, Lemma 4.6]).
 Lemma 2.3.
-Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $12\leq p_{1}\leq\tfrac{2c_{2}}{c_{3}^{2}}+$ 1 and for any $p_{2}\geq 2$ we have
-[…]
+Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $12\leq p_{1}\leq\tfrac{2c_{2}}{c_{3}^{2}}+$ 1 and for any $p_{2}\geq 2$, we have
+$\sup_{0\leq t<\infty}\mathbb{E}\big{[}|X_{t}|^{p_{1}}\big{]}<\infty\quad\text{and}\quad\sup_{0\leq t<\infty}\mathbb{E}\big{[}\big{|}\tfrac{1}{X_{t}}\big{|}^{p_{2}}\big{]}<\infty.$ (2.4)
 Based on the above moment integrability properties, one can also obtain the Hölder continuity of the solutions, as stated in the following lemmas quoted directly from [22].
 Lemma 2.4.
-[20, Lemma 4.4] Let conditions in Theorem 2.1 be satisfied with $\kappa+1>2\rho$. Then it holds that, for any $p\geq 1$ and $t,s\in[0,T]$,
+[22, Lemma 4.4] Let conditions in Theorem 2.1 be satisfied with $\kappa+1>2\rho$. Then it holds that, for any $p\geq 1$ and $t,s\in[0,T]$,
 $\|X_{t}-X_{s}\|_{L^{p}(\Omega;\mathbb{R})}$ $\leq C|t-s|^{\tfrac{1}{2}},$ (2.5)
@@
 Lemma 2.5.
-[20, Lemma 4.7] Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $t,s\in[0,T]$ it holds that
+[22, Lemma 4.7] Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $t,s\in[0,T]$, it holds that
 $\|X_{t}-X_{s}\|_{L^{q_{1}}(\Omega;\mathbb{R})}$ $\leq C|t-s|^{\tfrac{1}{2}},\quad 2\leq q_{1}\leq\tfrac{1}{\kappa}\big{(}\tfrac{2c_{2}}{c_{3}^{2}}+1\big{)},$ (2.7)
 $\|X_{t}^{-1}-X_{s}^{-1}\|_{L^{q_{2}}(\Omega;\mathbb{R})}$ $\leq C|t-s|^{\tfrac{1}{2}},\quad 2\leq q_{2}<\tfrac{1}{\kappa}\big{(}\tfrac{2c_{2}}{c_{3}^{2}}+1\big{)}.$ (2.8)
-Here and below, for notation simplicity the letter $C$ is slightly abused to denote a generic positive constant, which is independent of the time step-size and may vary for each appearance.
-As a straightforward consequence of the above lemmas, one can use the Hölder inequality to show the following estimates (consult [20, (4.21), (4.24)]).
+Here and below, for notation simplicity, the letter $C$ is slightly abused to denote a generic positive constant, which is independent of the time step-size and may vary for each appearance.
+As a straightforward consequence of the above lemmas, one can use the Hölder inequality to show the following estimates (consult [22, (4.21), (4.24)]).
 Lemma 2.6.
@@
     $\kappa+1=2\rho$, $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$,
-then for any $t,s\in[0,T]$ it holds that
+then for any $t,s\in[0,T]$, it holds that
 $\|f(X_{t})-f(X_{s})\|_{L^{2}(\Omega;\mathbb{R})}\leq C|t-s|^{\frac{1}{2}},\quad\|g(X_{t})-g(X_{s})\|_{L^{2}(\Omega;\mathbb{R})}\leq C|t-s|^{\frac{1}{2}},$ (2.9)
@@
 3 The explicit positivity-preserving Euler-type scheme
-In this section we aim to propose and analyze new strong approximation schemes for the generalized Aït-Sahalia model (2.2).
+In this section, we aim to propose and analyze new strong approximation schemes for the generalized Aït-Sahalia model (2.2).
 For $T\in(0,+\infty)$ and $N\in\mathbb{N}$, we construct a uniform mesh $\{t_{n}=nh\}_{n=0}^{N}$ on the time interval $[0,T]$ with a uniform step-size $h=\tfrac{T}{N}$.
@@
 Let $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$.
-Given any initial value $Y_{0}=X_{0}>0$, the scheme (3.1) admits unique, positive numerical solutions $\{Y_{n}\}_{n\in\mathbb{N}}$ for any stepsize $h=\tfrac{T}{N}>0$.
+Given any initial value $Y_{0}=X_{0}>0$, the scheme (3.1) admits unique, positive numerical solutions $\{Y_{n}\}_{n\in\mathbb{N}}$ for any step-size $h=\tfrac{T}{N}>0$.
 In other words, the scheme (3.1) is positivity preserving unconditionally.
-Proof: For $x>0$ we introduce the function
+Proof: For $x>0$, we introduce the function
 $G(x):=x-c_{-1}hx^{-1}.$ (3.3)
@@
 Assumption 3.3.
-Let $\kappa+1\geq 2\rho$. There exist a positive constant $L(\gamma)$, independent of the stepsize $h>0$, such that
+Let $\kappa+1\geq 2\rho$ with $\kappa,\rho>1$. There exists a positive constant $L(\gamma)$, independent of the step-size $h>0$, such that
 $\langle x,f_{h}(x)\rangle+\gamma|g_{h}(x)|^{2}\leq L<\infty,\quad\forall x\in(0,+\infty),$ (3.7)
-where $\gamma>0$ is assumed to be arbitrarily large for the case $\kappa+1>2\rho$ and $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ for the critical case $\kappa+1=2\rho$.
+where $\gamma\geq\tfrac{1}{2}$ is assumed to be arbitrarily large for the case $\kappa+1>2\rho$ and $\tfrac{1}{2}\leq\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ for the critical case $\kappa+1=2\rho$.
 Assumption 3.4.
-There exists some positive constant $m_{1},m_{2}\leq 2\gamma+1$ such that
+There exist some positive constants $m_{1},m_{2}\leq 2\gamma+1$ such that
 $\big{|}f(x)-f_{h}(x)\big{|}^{2}\leq Ch|x|^{m_{1}},\quad\big{|}g(x)-g_{h}(x)\big{|}^{2}\leq Ch|x|^{m_{2}},\quad\forall x\in(0,+\infty)$ (3.8)
@@
 where $\alpha$ is a positive constant satisfying $\alpha\geq\tfrac{1}{2}$.
-Here the choice of $f_{h}$ and $g_{h}$ are nothing but a kind of taming strategy, which was firstly proposed in [10] and later extensively studied by many authors, see, e.g., [19, 3, 16, 17, 12, 13].
+Here, the choice of $f_{h}$ and $g_{h}$ are nothing but a kind of taming strategy, which was firstly proposed in [11] and later extensively studied by many authors, see, e.g., [21, 3, 18, 19, 13, 14].
 To validate Assumption 3.2, noticing that $2\kappa\alpha\geq\kappa>\rho$, simple calculations infer that for any $x\in(0,+\infty)$
 $|f_{h}(x)|^{2\alpha}\leq\tfrac{c_{2}^{2\alpha}x^{2\kappa\alpha}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{2}^{2\alpha}h^{-\alpha},\quad|g_{h}(x)|^{\frac{2\kappa\alpha}{\rho}}\leq\tfrac{c_{3}^{\frac{2\kappa\alpha}{\rho}}x^{2\kappa\alpha}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{3}^{\frac{2\kappa\alpha}{\rho}}h^{-\alpha}.$ (3.10)
-Thus we have
-[…]
+Thus, we have
+$|f_{h}(x)|\leq c_{2}h^{-\frac{1}{2}}\quad\text{and}\quad|g_{h}(x)|\leq c_{3}h^{-\frac{\rho}{2\kappa}}\leq c_{3}T^{\frac{1}{2}-\frac{\rho}{2\kappa}}h^{-\frac{1}{2}},$ (3.11)
 which confirms Assumption 3.2.
@@
 $\big{|}f(x)-f_{h}(x)\big{|}$ $=\tfrac{c_{2}h^{\alpha}x^{(2\alpha+1)\kappa}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{2}h^{\alpha}x^{(2\alpha+1)\kappa},$ (3.14)
-$\big{|}g(x)-g_{h}(x)\big{|}$ $=\tfrac{c_{3}h^{\alpha}x^{(2\alpha+1)\rho}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{3}h^{\alpha}x^{(2\alpha+1)\rho}.$ (3.15)
-Let $m_{1}=2(2\alpha+1)\kappa$ and $m_{2}=2(2\alpha+1)\rho$. Since $\alpha\geq\tfrac{1}{2}$, (3.8) holds true on an additional condition $12(2\alpha+1)\kappa\leq 2\gamma+$ 1.
+$\big{|}g(x)-g_{h}(x)\big{|}$ $=\tfrac{c_{3}h^{\alpha}x^{{\color[rgb]{0,0,0}2\kappa\alpha+\rho}}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{3}h^{\alpha}x^{{\color[rgb]{0,0,0}2\kappa\alpha+\rho}}.$ (3.15)
+Let $m_{1}=2(2\alpha+1)\kappa$ and $m_{2}=2({\color[rgb]{0,0,0}2\kappa\alpha+\rho})$.
+Clearly, $m_{2}\leq 4\alpha\kappa+\kappa+1\leq m_{1}$.
+Since $\alpha\geq\tfrac{1}{2}$, (3.8) holds true on an additional condition $12(2\alpha+1)\kappa\leq 2\gamma+$ 1.
 In what follows we establish the integrability of the numerical approximations.
 Lemma 3.6.
-Let Assumptions 3.2, 3.3 hold. Then for any $12\leq p\leq 2\gamma+$ 1 and for any step-size $h>0$ it holds that
+Let Assumptions 3.2 and 3.3 hold. Then for any $12\leq p\leq 2\gamma+$ 1 and for any step-size $h>0$ it holds that
 $\sup_{0\leq n\leq N}\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}\leq C(1+|x_{0}|^{p})<\infty.$ (3.16)
@@
 $Z_{n+1}:=Y_{n+1}-c_{-1}hY_{n+1}^{-1}.$ (3.18)
-Next we introduce a time-continuous version of $Z_{n+1}$ in each time interval $[t_{n},t_{n+1}]$ as follows:
+Next, we introduce a time-continuous version of $Z_{n+1}$ in each time interval $[t_{n},t_{n+1}]$ as follows:
 $\widetilde{Z}_{t}^{(n)}:=Y_{n}+\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}(t-t_{n})+g_{h}(Y_{n})(W_{t}-W_{t_{n}}),\quad t\in[t_{n},t_{n+1}].$ (3.19)
@@
 Owing to (3.6), a rough moment bound can be derived for $\widetilde{Z}_{t}^{(n)}$.
-Indeed, by the Jensen inequality, for any $p\geq 1$ we have
+Indeed, by the Jensen inequality, for any $p\geq 1$, we have
 $\sup_{t_{n}\leq t\leq t_{n+1}}\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}\widetilde{Z}_{t}^{(n)}\big{|}^{2}\Big{)}^{p}\bigg{]}\leq\hat{C}\mathbb{E}\Big{[}\big{|}Y_{n}\big{|}^{2p}\Big{]}+\tilde{C}_{\varepsilon},$ (3.20)
-where $\varepsilon\geq 2c_{-1}h>0$ is a positive constant specified later and $\hat{C}$$,\tilde{C}_{\varepsilon}$ are positive constants independent of $h$. Thus
-$\mathbb{E}\bigg{[}\Big{(}\varepsilon+|Y_{n+1}|^{2}+c_{-1}^{2}h^{2}|Y_{n+1}|^{-2}-2c_{-1}h\Big{)}^{p}\bigg{]}=\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}Y_{n+1}-c_{-1}hY_{n+1}^{-1}\big{|}^{2}\Big{)}^{p}\bigg{]}\leq\hat{C}\mathbb{E}\Big{[}\big{|}Y_{n}\big{|}^{2p}\Big{]}+\tilde{C}_{\varepsilon}.$ (3.21)
-By noting $\varepsilon\geq 2c_{-1}h$, it holds that
+where $\varepsilon\geq 1+2c_{-1}h>1$ is a positive constant specified later and $\hat{C}$$,\tilde{C}_{\varepsilon}$ are positive constants independent of $h$.
+By noting $\varepsilon\geq 2c_{-1}h$, we infer
+$\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}\widetilde{Z}_{t_{n+1}}^{(n)}\big{|}^{2}\Big{)}^{p}\bigg{]}$ $=\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}Y_{n+1}-c_{-1}hY_{n+1}^{-1}\big{|}^{2}\Big{)}^{p}\bigg{]}$
+$=\mathbb{E}\bigg{[}\Big{(}\varepsilon+|Y_{n+1}|^{2}+c_{-1}^{2}h^{2}|Y_{n+1}|^{-2}-2c_{-1}h\Big{)}^{p}\bigg{]}$
+$\geq\mathbb{E}\big{[}\big{|}Y_{n+1}\big{|}^{2p}\big{]}.$ (3.21)
+Therefore, we deduce from (3.20) that
 $\mathbb{E}\big{[}\big{|}Y_{n+1}\big{|}^{2p}\big{]}\leq\hat{C}\mathbb{E}\big{[}\big{|}Y_{n}\big{|}^{2p}\big{]}+\tilde{C}_{\varepsilon}.$ (3.22)
-By iteration we get
-$\sup_{0\leq n\leq N}\mathbb{E}\big{[}\big{|}Y_{n}\big{|}^{2p}\big{]}<\infty,$ (3.23)
-which, together with (3.20), infers that
+By iteration, we get, for any $p\geq 1$ and $0\leq n\leq N$,
+$\mathbb{E}\big{[}\big{|}Y_{n}\big{|}^{2p}\big{]}\leq\hat{C}^{n}\mathbb{E}\big{[}\big{|}Y_{0}\big{|}^{2p}\big{]}+\tilde{C}_{\varepsilon}\sum_{i=0}^{n-1}\hat{C}^{i}<\infty,$ (3.23)
+which, together with (3.20), infers that, for any $p\geq 1$,
 $\sup_{0\leq n\leq N-1}\sup_{t_{n}\leq t\leq t_{n+1}}\mathbb{E}\big{[}\big{(}\varepsilon+\big{|}\widetilde{Z}_{t}^{(n)}\big{|}^{2}\big{)}^{p}\big{]}<\infty.$ (3.24)
@@
 Nevertheless, it suffices to allow the use of the Gronwall lemma later to arrive at uniform moment bounds.
-To do this, we let $\varepsilon\geq 1+2c_{-1}h$ in the following and denote
-$V_{\varepsilon,q}(x):=\big{(}\varepsilon+|x|^{2}\big{)}^{q},\quad q,x\in\mathbb{R}.$ (3.25)
-It can be easily checked that for any $q,q^{\prime}\in\mathbb{R}$,
+To do this, we denote
+$V_{\varepsilon,q}(x):=\big{(}\varepsilon+|x|^{2}\big{)}^{q},\quad q,x\in\mathbb{R},$ (3.25)
+where we recall $\varepsilon\geq 1+2c_{-1}h$. It can be easily checked that for any $q,q^{\prime}\in\mathbb{R}$,
 $V_{\varepsilon,q}(x)\cdot V_{\varepsilon,q^{\prime}}(x)=V_{\varepsilon,q+q^{\prime}}(x),\quad\big{(}V_{\varepsilon,q}(x)\big{)}^{q^{\prime}}=V_{\varepsilon,q\cdot q^{\prime}}(x).$ (3.26)
@@
 $V_{\varepsilon,q}^{\prime}(x)$ $=2qx\big{(}\varepsilon+|x|^{2}\big{)}^{q-1}=2qxV_{\varepsilon,q-1}(x),$ (3.28)
-$V_{\varepsilon,q}^{\prime\prime}(x)$ $=2q\big{(}\varepsilon+|x|^{2}\big{)}^{q-2}\big{[}\varepsilon+(2q-1)|x|^{2}\big{]}$ $\leq 2q\cdot\max{\{1,2q-1\}}\cdot V_{\varepsilon,q-1}(x).$
+$V_{\varepsilon,q}^{\prime\prime}(x)$ $=2q\big{(}\varepsilon+|x|^{2}\big{)}^{q-2}\big{[}\varepsilon+(2q-1)|x|^{2}\big{]}$ $\leq 2q\cdot\max{\{1,2q-1\}}\cdot V_{\varepsilon,q-1}(x)$
 and that
 $x\cdot V_{\varepsilon,q}(x)\leq\big{(}\varepsilon+|x|^{2}\big{)}^{\frac{1}{2}}\cdot V_{\varepsilon,q}(x)=V_{\varepsilon,\frac{1}{2}}(x)\cdot V_{\varepsilon,q}(x)=V_{\varepsilon,q+\frac{1}{2}}(x).$ (3.29)
-Equipped with these properties, we can start the estimation of $\widetilde{Z}_{t}^{(n)}$. The Itô formula together with (3.28) and (3.29) yields
+Equipped with these properties, we can start the estimation of $\widetilde{Z}_{t}^{(n)}$. The Itô formula together with (3.28) and (3.29) yields that for any $q\geq 1$,
 $V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})$ $=V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{t}\bigg{[}V_{\varepsilon,q}^{\prime}(\widetilde{Z}_{s}^{(n)})\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}+\tfrac{1}{2}V_{\varepsilon,q}^{\prime\prime}(\widetilde{Z}_{s}^{(n)})|g_{h}(Y_{n})|^{2}\bigg{]}\,\text{d}s$ (3.30)
@@
 $\quad+\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s},$
-where we also used (3.19) in the last step. In view of (3.7), for any $q\leq\gamma+\tfrac{1}{2}$, one deduces that
+where we also used (3.19) in the last step.
+In view of (3.7), for any $q\leq\gamma+\tfrac{1}{2}$, where $\gamma$ is assumed to be arbitrarily large for the case $\kappa+1>2\rho$ and $\tfrac{1}{2}\leq\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ for the critical case $\kappa+1=2\rho$, one deduces that
 $\big{\langle}Y_{n},\,f_{h}(Y_{n})\big{\rangle}+\tfrac{2q-1}{2}|g_{h}(Y_{n})|^{2}\leq L,$ (3.31)
@@
 $\quad+\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}.$
-Observing that
-$\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}\Bigg{]}=0,$ (3.35)
-taking expectations on both sides of (3.34) reveals that
+Recalling (3.19) and (3.25), one knows that the process $\widetilde{Z}_{s}^{(n)}$ is $\mathcal{F}_{s}$-measurable, $Y_{n}$ is $\mathcal{F}_{t_{n}}$-measurable by construction and both are integrable in view of (3.2) and (3.24).
+This ensures that the following expectation of the Itô stochastic integral vanishes:
+$\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}\Bigg{]}=0.$ (3.35)
+By observing that
+$|Y_{n}|^{2q}\leq\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}=V_{\varepsilon,q}(Y_{n})=V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)}),$ (3.36)
+we take expectations on both sides of (3.34) to show
 $\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})\Big{]}$ $\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q(c_{0}+c_{1}+1)-c_{1}-2)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,q}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch$ (3.37)
@@
 The case $q=1$ directly results in
-$\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t}^{(n)})\Big{]}$ $\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+c_{1}\cdot\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,1}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch.$ (3.37)
-Now we focus on the case $q>1$. Applying the Itô formula again on $V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})$ we get
+$\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t}^{(n)})\Big{]}$ $\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+{\color[rgb]{0,0,0}(2c_{0}+c_{1})}\cdot\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,1}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch.$ (3.38)
+Now, we focus on the case $q>1$. Applying the Itô formula again on $V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)}),$ we get
 $V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})$ $=V_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})+\int_{t_{n}}^{s}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})\cdot\big{(}-c_{0}+c_{1}Y_{n}\big{)}\,\text{d}r$ (3.39)
@@
 $\quad+\int_{t_{n}}^{s}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})g_{h}(Y_{n})\,\text{d}W_{r}.$
-Plugging (3.38) into (3.36) and noting that
-$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}=0,$ (3.39)
-we have
+Similar to (3.35), we note that
+$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}$ $=2q\int_{t_{n}}^{t}\mathbb{E}\bigg{[}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\bigg{]}\,\text{d}s$ (3.40)
+$=0.$
+Taking this into account and plugging (3.39) into (3.37), we have
 $\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t}^{(n)})\Big{]}$ (3.41)
@@
 $\quad+\underbrace{\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2q\int_{t_{n}}^{s}V_{\varepsilon,q-1}^{\prime}(\widetilde{Z}_{r}^{(n)})g_{h}(Y_{n})\,\text{d}W_{r}\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}.}_{=:D_{3}}$
-Using the Young inequality twice as well as the Jensen inequality on $D_{1}$ we get
+Utilizing the Young inequality of the form
+$a\cdot b\leq\tfrac{1}{2q}a^{2q}+\tfrac{2q-1}{2q}b^{\frac{2q}{2q-1}}=\tfrac{1}{2q}\Big{(}a^{2q}+(2q-1)b^{\frac{2q}{2q-1}}\Big{)},$
+we first derive
 $D_{1}$ $=4q(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\ \int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\cdot\big{(}-c_{0}+c_{1}Y_{n}\big{)}\,\text{d}r\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}$ (3.42)
-$\leq 2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\bigg{(}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}$
-$\quad+(2q-1)\Big{|}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\cdot\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{\frac{2q}{2q-1}}\bigg{)}\,\text{d}s\bigg{]}$
-$\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\bigg{(}(2q-2)^{2}\Big{|}h^{-1}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{q}{q-1}}$
-$\quad+(2q-2)\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\bigg{)}\,\text{d}s\bigg{]}$
-$\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{q+1},$
-where we used (3.6), (3.27), (3.29), the Hölder inequality and the moment inequality in the last step.
+$\leq 2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\,\text{d}s\bigg{]}$
+$\quad+2(q-1)(2q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{\frac{2q}{2q-1}}\,\text{d}s\bigg{]}.$
+Further, applying the Young inequality of another form
+$a\cdot b\leq\tfrac{2q-2}{2q-1}a^{\frac{2q-1}{2q-2}}+\tfrac{1}{2q-1}b^{2q-1}=\tfrac{1}{2q-1}\big{(}(2q-2)a^{\frac{2q-1}{2q-2}}+b^{2q-1}\big{)},$
+we get
+$D_{1}$ $\leq 2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\,\text{d}s\bigg{]}$ (3.43)
+$\quad+2(q-1)(2q-2)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h^{-1}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{2q}{2q-2}}\,\text{d}s\bigg{]}$
+$\quad+2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\,\text{d}s\bigg{]}.$
+For the first term, we can directly deduce that
+$|-c_{0}+c_{1}Y_{n}|^{2q}\leq C\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}=CV_{\varepsilon,q}(Y_{n})=CV_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)}).$
+For the second term, utilizing (3.27), (3.29) and the Hölder inequality, we get
+$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h^{-1}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{2q}{2q-2}}\,\text{d}s\bigg{]}$
+$\leq\int_{t_{n}}^{t}\mathbb{E}\bigg{[}\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{2q}{2q-2}}\bigg{]}\,\text{d}s$
+$\leq\int_{t_{n}}^{t}h^{\frac{q}{q-1}-1}\int_{t_{n}}^{s}\mathbb{E}\Big{[}\big{|}h^{-1}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{q}{q-1}}\Big{]}\text{d}r\text{d}s$
+$\leq\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s.$
+For the third term, the moment inequality [15, Theorem 7.1] together with (3.6) directly infers that
+$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\,\text{d}s\bigg{]}\leq Ch^{q+1}.$
+Thus, in summary, we have
+$D_{1}\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{q+1}.$ (3.44)
 Since the estimation of $D_{3}$ is similar and easy, we present it before treating $D_{2}$.
-By the generalized Itô isometry [8, Chapter 4, B], the Young inequality as well as (3.6), (3.27), (3.29), the Hölder inequality and the moment inequality we deduce that
+By the generalized Itô isometry [8, Chapter 4, B], the Young inequality as well as (3.6), (3.27), (3.29), the Hölder inequality, and the moment inequality [15, Theorem 7.1], we deduce that
 $D_{3}$ $=4q(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})f_{h}(Y_{n})|g_{h}(Y_{n})|^{2}\,\text{d}r\,\text{d}s\bigg{]}$ (3.45)
@@
 $=:D_{21}+D_{22}.$
-Employing the Young inequality, (3.27), (3.31) on $D_{21}$ as well as the Hölder inequality and the moment inequality we get
+Employing the Young inequality, (3.27), (3.31) on $D_{21}$ as well as the Hölder inequality and the moment inequality [15, Theorem 7.1], we get
 $D_{21}$ $\leq 4Lq(q-1)\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{(}\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}\cdot\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}\bigg{)}\,\text{d}s\Bigg{]}$ (3.47)
@@
 $\leq L(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{\frac{q}{2}+1}.$
-With regard to $D_{22}$, using the Young inequality twice again as well as the Hölder inequality, the moment inequality, (3.6), (3.19) and (3.27) on $D_{22}$ we deduce that
+With regard to $D_{22}$, using the Young inequality twice again as well as the Hölder inequality, the moment inequality, (3.6), (3.19), and (3.27) on $D_{22}$, we deduce that
 $D_{22}$ $\leq 2(q-1)\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}\bigg{(}(2q-1)\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})(\widetilde{Z}_{r}^{(n)}-Y_{n})\,\text{d}r\Big{|}^{\frac{2q}{2q-1}}$ (3.48)
@@
 $\leq\tfrac{2q-2}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}(2q-1)\mathbb{E}\bigg{[}\big{|}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{2q}{2q-1}}\cdot\big{|}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{|}^{\frac{2q}{2q-1}}\bigg{]}\,\text{d}r\text{d}s+Ch$
-$\leq\tfrac{2q-2}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\bigg{[}\big{|}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{2q-2}{2q-1}}\cdot\big{|}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{|}^{\frac{2q}{2q-1}}\bigg{]}\,\text{d}r\text{d}s+Ch$
+$\leq\tfrac{(2q-2){\color[rgb]{0,0,0}(2q-1)}}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\bigg{[}\big{|}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{2q-2}{2q-1}}\cdot\big{|}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{|}^{\frac{2q}{2q-1}}\bigg{]}\,\text{d}r\text{d}s+Ch$
 $\leq\tfrac{(2q-2)^{2}}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\,\text{d}r\text{d}s+Ch$
-$\quad+\tfrac{(2q-1)(2q-2)}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}+\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}\,\text{d}r\text{d}s$
-$\leq(2q-2)^{2}h\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+C\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch.$
+$\quad+\tfrac{{\color[rgb]{0,0,0}2q-2}}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}+\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}\,\text{d}r\text{d}s.$
+By the Jensen inequality, the Hölder inequality, the moment inequality [15, Theorem 7.1], and the fact that
+$\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\leq C\cdot\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}=C\cdot\big{(}\varepsilon+|\widetilde{Z}_{t_{n}}^{(n)}|^{2}\big{)}^{q}=C\cdot V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)}),$
+one can bound the integrand in last term of (3.48) as follows:
+$\mathbb{E}$ $\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}+\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}$
+$\leq 2^{2q-1}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}\bigg{|}^{2q}\Bigg{]}+2^{2q-1}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}$
+$\leq 2^{2q-1}(r-t_{n})^{2q-1}\int_{t_{n}}^{r}\mathbb{E}\Big{[}\big{|}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{|}^{2q}\Big{]}\,\text{d}r_{1}$
+$\quad+2^{2q-1}(r-t_{n})^{q-1}\int_{t_{n}}^{r}\mathbb{E}\Big{[}\big{|}g_{h}(Y_{n})\big{|}^{2q}\Big{]}\,\text{d}r_{1}$
+$\leq 2^{4q-2}(r-t_{n})^{2q}\mathbb{E}\Big{[}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\Big{]}+2^{4q-2}(r-t_{n})^{2q}\mathbb{E}\Big{[}\big{|}f_{h}(Y_{n})\big{|}^{2q}\Big{]}$
+$\quad+2^{2q-1}(r-t_{n})^{q}\mathbb{E}\Big{[}\big{|}g_{h}(Y_{n})\big{|}^{2q}\Big{]}$
+$\leq 2^{4q-2}(r-t_{n})^{2q}C\cdot\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+2^{4q-2}(r-t_{n})^{2q}h^{-q}+2^{2q-1}(r-t_{n})^{q}h^{-q}.$
+Plugging this into (3.48) shows
+${\color[rgb]{0,0,0}D_{22}\leq(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{2q+1}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+Ch}.$ (3.49)
 Combining all the above estimations and taking supremum reveal that
@@
 From now on, we set $\varepsilon=1+c_{-1}+2c_{-1}h$. On the one hand, one knows that
-$\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n+1}-c_{-1}hY_{n+1}^{-1}|^{2}\big{)}^{q}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}1+c_{-1}+2c_{-1}h+|Y_{n+1}|^{2}-2c_{-1}h+|Y_{n+1}|^{-2}\big{)}^{q}\Big{]}$ (3.49)
+$\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n+1}-c_{-1}hY_{n+1}^{-1}|^{2}\big{)}^{q}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}1+c_{-1}+2c_{-1}h+|Y_{n+1}|^{2}-2c_{-1}h+{\color[rgb]{0,0,0}c_{-1}^{2}h^{2}}|Y_{n+1}|^{-2}\big{)}^{q}\Big{]}$ (3.53)
 $\geq\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n+1}|^{2}\big{)}^{q}\Big{]}.$
-On the other hand,
+On the other hand, by noting
+$1+c_{-1}+2c_{-1}h+|Y_{n}|^{2}\leq 1+2h+c_{-1}(1+2h)+(1+2h)|Y_{n}|^{2}=(1+2h)\big{(}1+c_{-1}+|Y_{n}|^{2}\big{)},$
+one can deduce
 $\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}1+c_{-1}+2c_{-1}h+|Y_{n}|^{2}\big{)}^{q}\Big{]}$ (3.54)
@@
 $\leq\text{e}^{2hq}\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n}|^{2}\big{)}^{q}\Big{]}.$
-Thus we conclude that
+Thus, we conclude that
 $\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n+1}|^{2}\big{)}^{q}\Big{]}\leq\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n}|^{2}\big{)}^{q}\Big{]}\text{e}^{Ch}+Che^{Ch}.$ (3.55)
-By iteration we arrive at
+By iteration, we arrive at
 $\sup_{1\leq n\leq N}\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{n}|^{2})^{q}\Big{]}$ $\leq\sup_{1\leq n\leq N}\mathbb{E}\Big{[}\big{(}1+c_{-1}+|Y_{0}|^{2}\big{)}^{q}\Big{]}\text{e}^{Cnh}+C<\infty,$ (3.56)
@@
 Corollary 3.7.
-Let Assumptions 3.2, 3.3, 3.4 hold. Then it holds that
+Let Assumptions 3.2, 3.3, and 3.4 hold. Then it holds that
 $\mathbb{E}\Big{[}\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\Big{]}\leq Ch,\quad\mathbb{E}\Big{[}\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\Big{]}\leq Ch.$ (3.57)
@@
 4 Mean-square convergence rate of the proposed scheme
-In this section we present the analysis of the mean-square convergence rate for the proposed scheme on the mesh grid.
+In this section, we present the analysis of the mean-square convergence rate for the proposed scheme on the mesh grid.
 Two cases are considered including the non-critical case $\kappa+1>2\rho$ and the general critical case $\kappa+1=2\rho$.
@@
 Let $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{n}\}_{0\leq n\leq N}$ be defined by (2.2) and (3.1), respectively.
-Let Assumptions 3.2, 3.3 and 3.4 hold, with additionally one of the following conditions standing:
+Let Assumptions 3.2, 3.3, and 3.4 hold, with additionally one of the following conditions standing:
   1. (1)
@@
 $\quad+\delta|g(X_{t_{n}})-g(Y_{n})|^{2}|\Delta W_{n}|^{2}+\tfrac{1}{\delta}|R_{n+1}|^{2},$
-where $\delta>0$ is some positive constant. By noting that
-$\big{\langle}e_{n+1},X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{\rangle}=\int_{0}^{1}-\Big{(}Y_{n+1}+\theta(X_{t_{n+1}}-Y_{n+1})\Big{)}^{-2}\text{d}\theta\cdot(X_{t_{n+1}}-Y_{n+1})^{2}\leq 0$ (4.5)
-and that
-$\mathbb{E}\big{[}\langle e_{n},(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\rangle\big{]}=0,\,\mathbb{E}\big{[}\langle e_{n},R_{n+1}^{(2)}\rangle\big{]}$ $=0,$
-$\mathbb{E}\big{[}\big{\langle}f(X_{t_{n}})-f(Y_{n}),(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\big{\rangle}\big{]}$ $=0,$
-we thus take expectations on both sides of (4.4) and do some rearrangements to arrive at
+where $\delta>0$ is some positive constant. By noting that $e_{n}$, $f(X_{t_{n}})-f(Y_{n})$, $g(X_{t_{n}})-g(Y_{n})$ are $\mathcal{F}_{t_{n}}$-measurable, one can easily see that
+$\mathbb{E}\big{[}\langle e_{n},(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\rangle\big{]}$ $=0,$
+$\mathbb{E}\big{[}\big{\langle}f(X_{t_{n}})-f(Y_{n}),(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\big{\rangle}\big{]}$ $=0.$
+Since $e_{n}$ is $\mathcal{F}_{t_{n}}$-measurable, the property of stochastic integrals implies that
+$\mathbb{E}\big{[}\langle e_{n},R_{n+1}^{(2)}\rangle\big{]}=\mathbb{E}\bigg{[}\int_{t_{n}}^{t_{n+1}}e_{n}\Big{(}\big{(}g(X_{s})-g(X_{t_{n}})\big{)}+\big{(}g(Y_{n})-g_{h}(Y_{n})\big{)}\Big{)}\,\text{d}W_{s}\bigg{]}=0.$
+In addition, we note that
+$\big{\langle}e_{n+1},X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{\rangle}=\int_{0}^{1}-\Big{(}Y_{n+1}+\theta(X_{t_{n+1}}-Y_{n+1})\Big{)}^{-2}\text{d}\theta\cdot(X_{t_{n+1}}-Y_{n+1})^{2}\leq 0.$ (4.5)
+Therefore, we take expectations on both sides of (4.4) and do some rearrangements to arrive at
 $\mathbb{E}\big{[}|e_{n+1}|^{2}\big{]}$ $\leq(1+(1+3c_{1})h+(c_{1}^{2}+2c_{1})h^{2}+c_{1}^{2}h^{3})\mathbb{E}\big{[}|e_{n}|^{2}\big{]}+(2+c_{1}h)h^{2}\mathbb{E}\big{[}|f(X_{t_{n}})-f(Y_{n})|^{2}\big{]}$ (4.6)
@@
 $\leq(X_{t_{n}}-Y_{n})^{2}\int_{0}^{1}\big{[}(1+\delta)c_{3}^{2}\rho^{2}(\theta X_{t_{n}}+(1-\theta)Y_{n})^{2\rho-2}-2c_{2}\kappa(\theta X_{t_{n}}+(1-\theta)Y_{n})^{\kappa-1}\big{]}\,\text{d}\theta.$
-In view of this, we claim that
-$(1+\delta)h\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}+2h\mathbb{E}\big{[}\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle\big{]}\leq\tilde{L}(X_{t_{n}}-Y_{n})^{2}.$ (4.8)
-Indeed, for the case $\kappa+1>2\rho$, the assertion is obvious by noting $\kappa-1>2\rho-2$ and $02c_{2}\kappa>$ 0.
-For the critical case $\kappa+1=2\rho$, the assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ implies $\tfrac{c_{2}}{c_{3}^{2}}>\kappa/2$ and thus $\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-1>0$.
-As a result, one can choose $10<\delta\leq\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-$ 1 such that $(1+\delta)c_{3}^{2}\rho^{2}-2c_{2}\kappa\leq 0$, which ensures that (4.8) still remains valid.
+Here, we claim that for any $u\geq 0$,
+$(1+\delta)c_{3}^{2}\rho^{2}u^{2\rho-2}-2c_{2}\kappa u^{\kappa-1}\leq\tilde{L}.$ (4.8)
+Indeed, for the case $\kappa+1>2\rho$, the assertion is obvious by noting $\kappa-1>2\rho-2$ and the fact that the coefficient $-2c_{2}\kappa$ of the higher-order term $u^{\kappa-1}$ is negative.
+For the critical case $\kappa+1=2\rho$, the assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ implies that $\tfrac{c_{2}}{c_{3}^{2}}>\tfrac{\kappa}{2}+\tfrac{3}{2}(\kappa-1)>\tfrac{\kappa}{2}$ and thus $\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-1>\tfrac{\kappa^{2}}{\rho}-1>\rho-1>0$.
+As a result, one can choose $10<\delta\leq\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-$ 1 such that $(1+\delta)c_{3}^{2}\rho^{2}-2c_{2}\kappa\leq 0$, which ensures that the coefficient of $u^{\kappa-1}=u^{2\rho-2}$ is still negative.
+A combination of the two cases confirms (4.8).
+Bearing (4.8) in mind, one can readily derive from (4.7) that
+$(1+\delta)h\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}+2h\mathbb{E}\big{[}\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle\big{]}\leq\tilde{L}{\color[rgb]{0,0,0}h\mathbb{E}\big{[}|X_{t_{n}}-Y_{n}|^{2}\big{]}}.$ (4.9)
 Taking (4.9) into account, one can deduce from (4.6) that
-$\mathbb{E}\big{[}|e_{n+1}|^{2}\big{]}$ $\leq(1+(1+3c_{1})h+(c_{1}^{2}+2c_{1})h^{2}+c_{1}^{2}h^{3})\mathbb{E}\big{[}|e_{n}|^{2}\big{]}+(2+c_{1}h)h^{2}\mathbb{E}\big{[}|f(X_{t_{n}})-f(Y_{n})|^{2}\big{]}$ (4.9)
+$\mathbb{E}\big{[}|e_{n+1}|^{2}\big{]}$ $\leq(1+(1+3c_{1})h+(c_{1}^{2}+2c_{1})h^{2}+c_{1}^{2}h^{3}{\color[rgb]{0,0,0}+\tilde{L}})\mathbb{E}\big{[}|e_{n}|^{2}\big{]}+(2+c_{1}h)h^{2}\mathbb{E}\big{[}|f(X_{t_{n}})-f(Y_{n})|^{2}\big{]}$ (4.10)
 $\quad+(2+\tfrac{1}{\delta})\mathbb{E}\big{[}|R_{n+1}|^{2}\big{]}+\tfrac{1}{h}\mathbb{E}\big{[}|R_{n+1}^{(1)}|^{2}\big{]}.$
@@
     $\kappa+1>2\rho$.
-    Due to Lemma 2.2 and Lemma 3.6, one can infer that
+Noticing that $\gamma$ is an arbitrary large constant ensuring $\kappa\leq\gamma+\tfrac{1}{2}$ in the non-critical case due to Assumption 3.3, with the aid of Lemma 2.2 and Lemma 3.6, one can infer that
     $\mathbb{E}\big{[}\big{|}f(X_{t_{n}})-f(Y_{n})\big{|}^{2}\big{]}\leq C\quad\text{and}\quad\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}\leq C.$ (4.11)
@@
     $\mathbb{E}\big{[}\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch,\quad\mathbb{E}\big{[}\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch.$ (4.12)
-    Equipped with (4.11) as well as Lemmas 2.4, 2.6 one can directly obtain
-    $\mathbb{E}\big{[}\big{|}R_{n+1}^{(1)}\big{|}^{2}\big{]}\leq Ch^{3},\quad\mathbb{E}\big{[}\big{|}R_{n+1}^{(2)}\big{|}^{2}\big{]}\leq Ch^{2}.$ (4.12)
+    Equipped with (4.12) as well as (4.3), Lemmas 2.4 and 2.6 one can directly deduce by the Jensen inequality and the Hölder inequality that
+    $\mathbb{E}\Big{[}\big{|}R_{n+1}^{(1)}\big{|}^{2}\Big{]}$ $\leq 4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}c_{-1}(X_{s}^{-1}-X_{t_{n+1}}^{-1})\text{d}s\Big{|}^{2}\bigg{]}+4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}c_{1}(X_{s}-X_{t_{n}})\text{d}s\Big{|}^{2}\bigg{]}$
+    $\quad+4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}(f(X_{s})-f(X_{t_{n}}))\text{d}s\Big{|}^{2}\bigg{]}+4\mathbb{E}\bigg{[}\Big{|}\int_{t_{n}}^{t_{n+1}}(f(Y_{n})-f_{h}(Y_{n}))\text{d}s\Big{|}^{2}\bigg{]}$
+    $\leq 4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}c_{-1}(X_{s}^{-1}-X_{t_{n+1}}^{-1})\big{|}^{2}\Big{]}\text{d}s+4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}c_{1}(X_{s}-X_{t_{n}})\big{|}^{2}\Big{]}\text{d}s$
+    $\quad+4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}(f(X_{s})-f(X_{t_{n}}))\big{|}^{2}\Big{]}\text{d}s+4h\int_{t_{n}}^{t_{n+1}}\mathbb{E}\Big{[}\big{|}(f(Y_{n})-f_{h}(Y_{n}))\big{|}^{2}\Big{]}\text{d}s$
+    $\leq Ch^{3}.$
+    In a similar way, but with the Itô isometry leads to
+    $\mathbb{E}\big{[}\big{|}R_{n+1}^{(2)}\big{|}^{2}\big{]}\leq Ch^{2}.$ (4.13)
   2. Case 2:
     $\kappa+1=2\rho.$
-    The difference for the critical case comes from the extra restriction that $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$, as required by Assumption 3.3, which together with Lemmas 2.3, 3.6 ensures the moment boundedness of $\mathbb{E}\big{[}|X_{t}|^{p}\big{]}$ and $\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}$ for $p\leq 2\gamma+1\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$.
+    The difference for the critical case comes from the extra restriction that $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$, as required by Assumption 3.3, which together with Lemmas 2.3 and 3.6 ensures the moment boundedness of $\mathbb{E}\big{[}|X_{t}|^{p}\big{]}$ and $\mathbb{E}\big{[}|Y_{n}|^{p}\big{]}$ for $p\leq 2\gamma+1\leq\tfrac{2c_{2}}{c_{3}^{2}}+1$.
 Therefore, on the condition $\tfrac{c_{2}}{c_{3}^{2}}>\kappa-\tfrac{1}{2}$, one can derive $2\kappa$-th moment bounds for both analytic and numerical solutions
     $\mathbb{E}\big{[}\big{|}f(X_{t_{n}})-f(Y_{n})\big{|}^{2}\big{]}\leq C\quad\text{and}\quad\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}\leq C.$ (4.14)
-    Recalling (3.8) one knows that
+    Recalling (3.8), one knows that
     $\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\leq Ch|Y_{n}|^{m_{1}},\quad\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\leq Ch|Y_{n}|^{m_{2}},$ (4.15)
@@
     $\mathbb{E}\big{[}\big{|}f(Y_{n})-f_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch,\quad\mathbb{E}\big{[}\big{|}g(Y_{n})-g_{h}(Y_{n})\big{|}^{2}\big{]}\leq Ch,$ (4.16)
-    by Corollary 3.7. This together with Lemma 2.5 and 2.6 helps us to deduce that, on the more strict assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$,
+    by Corollary 3.7. This together with Lemmas 2.5 and 2.6 similarly helps us to deduce that, on the more strict assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$,
     $\mathbb{E}\big{[}\big{|}R_{n+1}^{(1)}\big{|}^{2}\big{]}\leq Ch^{3},\quad\mathbb{E}\big{[}\big{|}R_{n+1}^{(2)}\big{|}^{2}\big{]}\leq Ch^{2}.$ (4.17)
@@
 Let $f_{h}$ and $g_{h}$ be defined as in Example 3.5.
-Suppose that Assumptions 3.2, 3.3 and 3.4 hold, with additionally one of the following conditions standing:
+Suppose that Assumptions 3.2, 3.3, and 3.4 hold, with additionally one of the following conditions standing:
   1. (1)
@@
     $\kappa+1=2\rho$, $\tfrac{c_{2}}{c_{3}^{2}}\geq(2\alpha+1)\kappa-\tfrac{1}{2}$.
-Then there exists a constant $C>0$ independent of the step-size $h$ such that
+Then, there exists a constant $C>0$ independent of the step-size $h$ such that
 $\sup_{0\leq n\leq N}\mathbb{E}\big{[}\big{|}X_{t_{n}}-Y_{n}\big{|}^{2}\big{]}\leq Ch.$ (4.20)
-We would like to mention that, for the general critical case $\kappa+1=2\rho$ one can observe a slightly more strict restriction on the model parameters than that for the implicit methods (STMs) in [20], where $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ is required.
+We would like to mention that, for the general critical case $\kappa+1=2\rho$, one can observe a slightly more strict restriction on the model parameters than that for the implicit methods (STMs) in [22], where $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ is required.
 This is caused by the need of $m_{1},m_{2}\leq 2\gamma+1$ in order to confirm Assumption 3.4.
-More precisely, as indicated by Example 3.5, for the general critical case, $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ as well as $12(2\alpha+1)\rho=m_{2}<m_{1}=2(2\alpha+1)\kappa\leq 2\gamma+$ 1 are required, leading to $(2\alpha+1)\kappa\leq\tfrac{c_{2}}{c_{3}^{2}}+\tfrac{1}{2}$ as mentioned.
+More precisely, as indicated by Example 3.5, for the general critical case, $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ as well as $12{\color[rgb]{0,0,0}(2\alpha\kappa+\rho)}=m_{2}<m_{1}=2(2\alpha+1)\kappa\leq 2\gamma+$ 1 are required, leading to $(2\alpha+1)\kappa\leq\tfrac{c_{2}}{c_{3}^{2}}+\tfrac{1}{2}$ as mentioned.
 5 Numerical experiments
@@
 $\mathrm{d}X_{t}=$ $(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\mathrm{d}t+c_{3}X_{t}^{\rho}\mathrm{d}W_{t},\quad X_{0}=1>0.$ (5.1)
-The semi-implicit tamed Euler-type method (TEM) in Example 3.5 with $\alpha=\tfrac{1}{2}$ is applied to approximate the the model:
+The semi-implicit tamed Euler-type method (TEM) in Example 3.5 with $\alpha=\tfrac{1}{2}$ applied to the model reads:
 $Y_{n+1}=Y_{n}+c_{-1}Y_{n+1}^{-1}h+\big{(}-c_{0}+c_{1}Y_{n}-\tfrac{c_{2}Y_{n}^{\kappa}}{1+\sqrt{h}Y_{n}^{\kappa}}\big{)}h+\tfrac{c_{3}Y_{n}^{\rho}}{1+\sqrt{h}Y_{n}^{\kappa}}\Delta W_{n}.$ (5.2)
-The following two sets of parameters are taken as numerical tests.
+The following three sets of parameters are taken as numerical tests.
 Example 5.1.
@@
 Example 5.2.
-$\kappa=2,\rho=1.5,c_{-1}=1.5,c_{0}=2,c_{1}=1,c_{2}=4,c_{3}=1$.
-Note that Example 5.1 satisfies the non-critical condition $\kappa+1>2\rho$, while Example 5.2 satisfies the general critical condition $\kappa+1=2\rho$.
-The strong convergence rates are tested with fixed total time $T=1$.
-The numerical approximations with the step-size $h_{exact}=2^{-14}$ is used to simulate the ”exact” solutions, while $h=2^{-i},i=4,5,...,9$ are used for numerical solutions over $10^{4}$ Brownian paths.
+${\color[rgb]{0,0,0}\kappa=3,\rho=2},c_{-1}=1.5,c_{0}=2,c_{1}=1,c_{2}=4,{\color[rgb]{0,0,0}c_{3}=0.5}$.
+Example 5.3.
+$\kappa=2,\rho=1.5,c_{-1}=2,c_{0}=3,c_{1}=4,c_{2}=7,c_{3}=\sqrt{2}$.
+Note that Example 5.1 satisfies the non-critical condition $\kappa+1>2\rho$, Example 5.2 satisfies the critical condition $\kappa+1=2\rho$, and Example 5.3, borrowed from [10, SET II], satisfies the critical condition with $\kappa=2$, $\rho=1.5$ and $\tfrac{c_{2}}{c_{3}^{2}}=2\kappa-\tfrac{1}{2}$.
+The strong convergence rates are tested with the endpoint $T=1$.
+For comparison, the backward Euler method (BEM) [20], which has been proved to be strong convergent with order 0.5 for the model [22], is also tested.
+Moreover, numerical approximations produced by the BEM with the step-size $h_{exact}=2^{-14}$ is used to simulate the “exact” solutions, while $h=2^{-i},i=4,5,...,9$ are used for numerical solutions over $10^{4}$ Brownian paths.
 The estimation error is calculated in terms of $e_{h}:=\bigg{(}\sup\limits_{0\leq n\leq N}\mathbb{E}\Big{[}|X_{t_{n}}-Y_{n}|^{2}\Big{]}\bigg{)}^{\tfrac{1}{2}}$.
-The stochastic theta method (STM) [20] with $\theta=1$, which coincides with the backward Euler-Maruyama method (BEM) [18], is also tested in comparison with TEM.
 Fig.
-1 plots the average sample error against various step-sizes on a log-log scale, with BEM in a black solid line and TEM in a red solid line.
-It can be easily observed the order $0.5$ strong convergence for both two schemes.
-By a least-squares fitting, the resulting convergence rate $q$ and the least squares residual are shown in Table 1, where one can also confirm the predicted convergence rate.
-Moreover, the computational costs of BEM and the proposed TEM scheme over $10^{4}$ Brownian paths are shown in Table 2 and 50 out of the $10^{4}$ paths are shown in Fig.
-2.
-It is easy to see that the TEM can preserve positivity and reduce computational costs sigficantly, compared with BEM, especially as the parameter $\kappa$ increases.
-Figure 1: Strong convergence rate of the semi-implicit taming Euler method under Example 5.1 (left) and Example 5.2 (right).
-Figure 2: 50 paths of numerical solutions for BEM and TEM
-Table 1: A least squares fit for the convergence rate $q$
+1 and Fig.
+2 plot the average sample errors against various step-sizes on a log-log scale for Examples 5.1, 5.2, and 5.3, with BEM in a blue solid line, TEM in a red solid line, and the boundary preserving explicit scheme (BPES) proposed in [10] in a green solid line.
+The expected order $0.5$ of strong convergence is easily detected for both BEM and TEM.
+By a least square fitting, the resulting convergence rate $q$ and the least squares residual are shown in Table 1, where one can also confirm the predicted convergence rate.
+For Example 5.3 of the particular critical case $\kappa=2,\rho=1.5$, the BPES shows order one convergence and outperforms the other two order one-half schemes in terms of accuracy.
+Nevertheless, the BPES is simply designed for the particular critical case $\kappa=2,\rho=1.5$ (see [10] for more details), but the proposed explicit scheme, TEM, works for all cases.
+In addition, the computational costs of BEM and the proposed TEM scheme over $10^{4}$ Brownian paths are shown in Table 2, and 50 out of the $10^{4}$ paths are shown in Fig.
+3.
+It is easy to see that the TEM can preserve positivity and reduce computational costs significantly, compared with BEM, especially as the parameter $\kappa$ increases.
+Figure 1: Strong convergence rate of TEM and BEM for Example 5.1 (left) and Example 5.2 (right).
+Figure 2: Strong convergence rate of TEM and BEM for Example 5.3.
+Figure 3: Fifty paths of numerical solutions for BEM and TEM
+Table 1: A least square fit for the convergence rate $q$
       BEM       TEM
-Case 1 $q=0.5409$, $resid=0.0269$ $q=0.5173$, $resid=0.0128$
-Case 2 $q=0.5558$, $resid=0.0698$ $q=0.5288$, $resid=0.0326$
+      Eg. 5.1       $q=0.5409$, $resid=0.0269$        $q=0.4458$, $resid=0.0074$
+      Eg. 5.2        $q=0.6658$, $resid=0.1161$       $q=0.5347$, $resid=0.0524$
+      Eg. 5.3       $q=0.5095$, $resid=0.0352$       $q=0.4906$, $resid=0.0174$
 Table 2: Time cost (seconds) for BEM and TEM over $10^{4}$ Brownian paths
                  BEM                  TEM
-Example 5.1 99.937 0.193
-Example 5.2 47.997 0.094
+                 Example 5.1                  135.586                  51.905
+                 Example 5.2                  78.977                  36.189
+                 Example 5.3                  34.085                  25.358
 6 Conclusion
-In this paper we propose and analyze an explicit time-stepping scheme for the strong approximations of the generalized Aït-Sahalia model.
+In this paper, we propose and analyze an explicit time-stepping scheme for the strong approximations of the generalized Aït-Sahalia model.
 Based on a semi-implicit strategy and proper modifications, the numerical scheme is designed to be positivity preserving and easily implementable with cheap computational costs compared with existing positivity preserving schemes.